### v1

# Quaternion Multi-focus Color Image Fusion

Abstract
Multi-focus color image fusion refers to integrating multiple partially focused color images to create a single all-in-focus color image. However, existing methods struggle with complex real-world scenarios due to limitations in handling color information and intricate textures. To address these challenges, this paper proposes a quaternion multi-focus color image fusion framework to perform high-quality color image fusion completely in the quaternion domain. This framework introduces 1) a quaternion sparse decomposition model to jointly learn fine-scale image details and structure information of color images in an iterative fashion for high-precision focus detection, 2) a quaternion base-detail fusion strategy to individually fuse base-scale and detail-scale results across multiple color images for preserving structure and detail information, and 3) a quaternion structural similarity refinement strategy to adaptively select optimal patches from initial fusion results and obtain the final fused result for preserving fine details and ensuring spatially consistent outputs. Extensive experiments demonstrate that the proposed framework outperforms state-of-the-art methods.
Index Terms:
Multi-focus color image fusion, quaternion color image fusion, quaternion sparse decomposition.
publicationid: pubid: $0$–$00$ 2025 IEEE
I Introduction
A single color image often fails to capture all objects in focus due to optical limitations of cameras [1]. When imaging scenes with objects at varying distances, only regions within the focal plane are focused while areas outside this plane exhibit defocus blur. To circumvent this limitation, multi-focus color image fusion (MCIF) techniques integrate multiple partially focused color images of the same scene using focus detection and adaptive fusion strategies to obtain an all-in-focus color image [2]. This technology has numerous applications across diverse domains including digital photography, preprocessing operations [3], and target recognition [4, 5].
Existing MCIF methods can be broadly categorized into deep learning-based and conventional methods. Currently, deep learning-based methods have extensively promoted the development of MCIF to enhance the ability of learning multi-scale features to distinguish focused pixels [6, 7]. Since it is impractical to obtain perfect ground truth (i.e., all-in-focus color images) for real-world scenes [8], many deep learning-based works heavily rely on synthetic training data with a large-scale training process. Unlike deep learning-based MCIF methods, conventional methods only depend on human prior knowledge to distinguish focused objects without the training phase [9].
Existing conventional MCIF methods often treat color image fusion as a direct extension of multi-focus grayscale image fusion techniques by channel-wise processing [10] or grayscale conversion [11]. For channel-wise processing, these MCIF methods simply treat color channels as three independent grayscale images and directly apply image fusion techniques to each channel to obtain the fused color image. For instance, Yang et al. [10] extended sparse coding techniques to MCIF by separately measuring focus levels in each channel and fusing sparse coefficients channel-by-channel. This may introduce spectral-spatial inconsistencies and color distortions in the fused color images [12, 13].
For grayscale conversion, existing conventional MCIF methods first convert the input color images into grayscale images and perform intensity-driven focus detection [14, 11] to derive focus maps from grayscale inputs to guide the fusion of original color channels [15]. For instance, [16] averaged color channels into a grayscale image and obtained focus maps by multi-scale morphological filtering. [17] transformed the color image into the YUV color space and detected focus levels employing the zero crossing and canny edge detectors. Although these methods avoid explicit channel separation, they fail to detect multi-channel sharpness variations among the input color images and may produce inaccurate focus maps under complex real-world scenarios. This leads to ghosting artifacts or blurred boundaries in the fused results [15, 13]. Therefore, it remains a challenge to perform high-quality color image fusion in real-world applications. The limitations of conventional MCIF methods lie mainly in two key challenges: (1) small-scale focus detection failure: existing techniques struggle to distinguish and retain small focused regions within images containing extensive defocused areas based on grayscale focus detection [4]; (2) boundary artifacts: focus maps derived from intensity-based or edge-based focus detection often produce inaccurate results with visual artifacts introduced particularly in uncertain boundary regions (between focused and defocused areas) and low-gradient regions due to defocus spread effects [15, 18].
Recently, a quaternion higher-order singular value decomposition (QHOSVD) model [19] was developed to perform quaternion-domain multi-focus color image fusion by fully considering highly-correlated color information. QHOSVD treats three color channels as a unified entity of three imaginary parts in a pure quaternion matrix using quaternion representation [20, 21] and extends this representation into a third-order quaternion tensor. This effectively preserves cross-channel relationships during color image processing [22, 23, 24]. Taking advantage of higher-order singular value decomposition in the quaternion domain, QHOSVD extracted focus levels corresponding to each source color image and employed a simple patch-wise fusion strategy to ensure spatial consistency and color fidelity [19]. However, QHOSVD still fails in focus information preservation and artifact suppression in complex real-world scenarios due to defocus spread effects [18].
To address these issues, we propose a quaternion multi-focus color image fusion (QMCIF) framework to produce high-quality fused color images. Our main contributions are presented as follows:
  • We propose a QMCIF framework to perform high-quality color image fusion completely in the quaternion domain. QMCIF is flexibly extended from fusing only two color images at a time into fusing multiple color images simultaneously under various challenging scenarios.
  • To effectively measure the focus level in both small focused areas and uncertain boundary regions, QMCIF introduces a quaternion sparse decomposition model to jointly learn fine-scale image details and structure information of color images in an iterative fashion.
  • To preserve color structure and detail information, QMCIF introduces a patch-wise quaternion base-detail fusion strategy to fuse multiple partially focused color images in detail-scale and base-scale individually.
  • To balance the trade-off between focus information preservation and artifact removal, QMCIF further introduces a quaternion structural similarity refinement strategy to flexibly form the final fused results.
  • Extensive experiments on various datasets demonstrate that QMCIF outperforms the state-of-the-art methods.
The rest of this paper is organized as follows: Section II presents the preliminaries of quaternion algebra. Section III introduces our framework and extends our framework to multi-color image fusion. Section IV presents the experiments and comparisons. Finally, Section V gives the conclusions.
II Preliminaries
This section presents quaternion representations of scalar numbers, vectors and matrices in detail. Table I shows the main notations and mathematic symbols used in this paper.
TABLE I: Notations and mathematical symbols.
$\mathbbm{R}$, $\mathbbm{H}$ real space, quaternion space
$a$, $\bm{\mathrm{a}}$, $\bm{\mathrm{A}}$ real scalar number, vector, matrix
$\dot{a}$, $\dot{\bm{\mathrm{a}}}$, $\dot{\bm{\mathrm{A}}}$ quaternion scalar number, vector, matrix
${\bm{\mathrm{I}}}_{d}$, $\dot{\bm{\mathrm{I}}}_{d}$ real identity matrix, quaternion identity matrix
${(\cdot)}^{\mathrm{T}}$$\overline{(\cdot)}$,${(\cdot)}^{\mathrm{H}}$, ${(\cdot)}^{-1}$ transpose, conjugate , conjugate transpose and inverse representation
${\|\cdot\|}_{1}$, ${\|\cdot\|}_{F}$, ${\|\cdot\|}_{*}$ ${\ell}_{1}$ norm, Frobenius norm, nuclear norm
The set of quaternions $\mathbbm{H}$ defines a 4-components normed algebra [25] over the real numbers $\mathbbm{R}$ (i.e., basis $\{1,i,j,k\}$) as follows:
$\begin{split}\dot{q}=q_{a}+q_{b}i+q_{c}j+q_{d}k,\end{split}$
where $q_{a},q_{b},q_{c},q_{d}\in\mathbbm{R}$ are components of $\dot{q}$. $i,j,k$ are the imaginary parts such that
$\begin{split}i^{2}=j^{2}=k^{2}=ijk=-1,ij=-ji,ij=k\end{split}$
These relations imply that quaternion multiplication is non-commutative. For $\dot{q},\dot{p}\in\mathbbm{H}$, $\dot{q}\dot{p}\neq\dot{p}\dot{q}$.
A quaternion $\dot{q}$ is a pure quaternion number if its real part $q_{a}=0$, namely $\dot{q}=q_{b}i+q_{c}j+q_{d}k$. Next, we will present several definitions that will be used in this paper.
Definition 1.
Given a quaternion vector $\dot{\bm{\mathrm{q}}}=(\dot{q}_{s})\in\mathbbm{H}^{M}$, and a quaternion matrix $\dot{\bm{\mathrm{A}}}=({\dot{a}}_{s,t})\in\mathbbm{H}^{M\times N}$, where $s=1,\cdots,M$ and $t=1,\cdots,N$ are the row and column indices respectively.
  1. 1)
    conjugate transpose: $\dot{{\bm{\mathrm{A}}}}^{\mathrm{H}}=(\overline{\dot{a}_{t,s}})\in\mathbbm{H}^{N\times M}$
  2. 2)
    ${\ell}_{1}$ norm: ${\|\dot{\bm{\mathrm{q}}}\|}_{1}=\sum_{s=1}^{M}{|\dot{q}_{s}|}$
  3. 3)
    ${\ell}_{2}$ norm: ${\|\dot{\bm{\mathrm{q}}}\|}_{2}={(\sum_{s=1}^{M}{|\dot{q}_{s}|}^{2})}^{\frac{1}{2}}$
  4. 4)
    $Frobenius$ norm: ${\|\dot{\bm{\mathrm{A}}}\|}_{F}={(\sum_{s=1}^{M}\sum_{t=1}^{N}{{|\dot{a}_{s,t}|}^{2}})}^{\frac{1}{2}}$
Definition 2.
The rank of a quaternion matrix $\dot{\bm{\mathrm{A}}}$ is r if and only if $\dot{\bm{\mathrm{A}}}$ has r nonzero singular values [26].
Definition 3.
Given a color image $\bm{\mathrm{I}}$ in real three-dimensional space, its quaternion representation is defined as:
$\begin{split}\dot{\bm{\mathrm{I}}}(s,t)=\bm{\mathrm{I}}_{r}(s,t)i+\bm{\mathrm{I}}_{g}(s,t)j+\bm{\mathrm{I}}_{b}(s,t)k\end{split}$
where $\dot{\bm{\mathrm{I}}}(s,t)$ is the quaternion representation of the color image pixel at the location of $(s,t)$. $\bm{\mathrm{I}}_{r}$, $\bm{\mathrm{I}}_{g}$ and $\bm{\mathrm{I}}_{b}$ is red, green and blue channels in $\bm{\mathrm{I}}$ respectively.
Definition 4.
(Quaternion derivatives [27]) The quaternion derivatives of the real scalar function $f$ : $\mathbbm{H}^{M\times N}\rightarrow\mathbbm{R}$ with respect to $\dot{\bm{\mathrm{Q}}}$ $\in\mathbbm{H}^{M\times N}$ are defined by
$\frac{\partial f}{\partial\dot{\bm{\mathrm{Q}}}}=\begin{pmatrix}\frac{\partial f}{\partial{\dot{q}}_{1,1}}&\cdots&\frac{\partial f}{\partial{\dot{q}}_{1,N}}\\ \vdots&\ddots&\vdots\\ \frac{\partial f}{\partial{\dot{q}}_{M,1}}&\cdots&\frac{\partial f}{\partial{\dot{q}}_{M,N}}.\end{pmatrix}$
where $\frac{\partial f}{\partial{\dot{q}}}=\frac{1}{4}(\frac{\partial f}{\partial{q_{a}}}-\frac{\partial f}{\partial{q_{b}}}i-\frac{\partial f}{\partial{q_{c}}}j-\frac{\partial f}{\partial{q_{d}}}k)$. The quaternion derivation has the following properties. They make the quaternion derivation significantly different from real-valued and complex-valued ones.
  1. 1.
    Non-commutativity: $\frac{\partial f}{\partial q_{b}}$, $\frac{\partial f}{\partial q_{c}}$, $\frac{\partial f}{\partial q_{d}}$ cannot be swapped with $i$, $j$, $k$.
  2. 2.
    $\frac{\partial{\dot{q}}}{\partial{\dot{q}}}=\frac{\partial\overline{\dot{q}}}{\partial\overline{\dot{q}}}=1$, $\frac{\partial{\dot{q}}}{\partial\overline{\dot{q}}}=\frac{\partial\overline{\dot{q}}}{\partial{\dot{q}}}=-\frac{1}{2}$.
  3. 3.
    Product rule: $\frac{\partial fg}{\partial{\dot{q}}}=f\frac{\partial g}{\partial{\dot{q}}}+\frac{\partial f}{\partial{\dot{q}}^{f}}g\neq f\frac{\partial g}{\partial{\dot{q}}}+\frac{\partial f}{\partial{\dot{q}}}g$.
  4. 4.
    Chain rule: $\frac{\partial f(g)}{\partial{\dot{q}}}=\frac{\partial f}{\partial g}\frac{\partial g}{\partial{\dot{q}}}+\frac{\partial f}{\partial g^{\bm{i}}}\frac{\partial g^{\bm{i}}}{\partial{\dot{q}}}+\frac{\partial f}{\partial g^{\bm{j}}}\frac{\partial g^{\bm{j}}}{\partial{\dot{q}}}+\frac{\partial f}{\partial g^{\bm{k}}}\frac{\partial g^{\bm{k}}}{\partial{\dot{q}}}$.
Definition 5.
Quaternion Structural Similarity (QSSIM) [28] measures the similarity between two color images in the quaternion domain. It is defined by
$QSSIM({\dot{\bm{\mathrm{X}}}},{\dot{\bm{\mathrm{Y}}}})=\overline{(\dot{a})}\dot{b},$ (1)
where
$\dot{a}=\frac{2\overline{\dot{\mu}_{\dot{\bm{\mathrm{X}}}}}\dot{\mu}_{\dot{\bm{\mathrm{Y}}}}+C_{1}}{\overline{{\dot{\bm{\mathrm{\mu}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}+C_{1}},$ $\dot{b}=\frac{2\overline{\dot{\sigma}_{\dot{\bm{\mathrm{X}}}}}\dot{\sigma}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}{\overline{{\dot{\bm{\mathrm{\sigma}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}.$ (2)
$\dot{\bm{\mathrm{X}}}$ and $\dot{\bm{\mathrm{Y}}}$ are input patches. $\dot{\mu}_{\dot{\bm{\mathrm{X}}}}$ and $\dot{\sigma}_{\dot{\bm{\mathrm{X}}}}$ are denoted as mean and variance values of the image patch $\dot{\bm{\mathrm{X}}}$. Constants $C_{1}$ and $C_{2}$ are set as small as possible.
Figure 1: Flowchart of quaternion multi-focus color image fusion (QMCIF) framework.
Lemma 1.
(Quaternion nuclear norm [26]) For any $\lambda\geq 0$, quaternion matrix $\dot{\bm{\mathrm{Y}}}$ and $\dot{\bm{\mathrm{X}}}$ $\in\mathbbm{H}^{M\times N}$ both with the rank of $r$, the quaternion nuclear norm can be defined as
$\mathop{\arg\min}_{\dot{\bm{\mathrm{X}}}}\ \ \frac{1}{2}{\|\dot{\bm{\mathrm{Y}}}-\dot{\bm{\mathrm{X}}}\|}_{F}^{2}+\lambda{\|\dot{\bm{\mathrm{X}}}\|}_{*}.$
Its closed form solution is $\dot{\bm{\mathrm{U}}}S_{\lambda}(\Sigma)\dot{\bm{\mathrm{V}}}^{H}$ where $\dot{\bm{\mathrm{U}}}$ $\in\mathbbm{H}^{M\times r}$ and $\dot{\bm{\mathrm{V}}}$ $\in\mathbbm{H}^{r\times N}$ are the orthogonal quaternion matrices, $\Sigma$ is a diagonal real matrix with singular values and $S_{\lambda}(\Sigma)$ denotes the soft thresholding operator with parameter $\lambda$.
Lemma 2.
(Quaternion soft-thresholding operator [24]) Given $\dot{\bm{\mathrm{X}}}and\dot{\bm{\mathrm{Y}}}\in\mathbbm{H}^{M\times N}$, we solve the quaternion ${\ell}_{1}$ norm optimization problem with respect to $\dot{\bm{\mathrm{X}}}$ using a quaternion soft-thresholding operator.
$\dot{\bm{\mathrm{X}}}=\mathop{\arg\min}_{\dot{\bm{\mathrm{X}}}}\ \ \tau{\|\dot{\bm{\mathrm{X}}}\|}_{1}+\frac{1}{2}\|\dot{\bm{\mathrm{X}}}-\dot{\bm{\mathrm{Y}}}\|{{}_{F}^{2}},$
$\dot{\bm{\mathrm{X}}}(:,i)=\left\{\begin{aligned} &\frac{{\|\dot{\bm{\mathrm{Y}}}(:,i)\|}_{1}-\tau}{{\|\dot{\bm{\mathrm{Y}}}(:,i)\|}_{1}}\dot{\bm{\mathrm{Y}}}(:,i),&&{\|\dot{\bm{\mathrm{Y}}}(:,i)\|}_{1}>\tau\\ &0,&&{\rm otherwise}\end{aligned}\right.$
III Proposed Framework
This section introduces our quaternion multi-focus color image fusion (QMCIF) framework. Section A provides an overview of our QMCIF framework. Section B introduces our quaternion sparse decomposition model and its optimization in detail. Section C presents our quaternion base-detail fusion strategy. Section D presents our quaternion structural similarity refinement strategy. Section E extends our framework to multiple-color image fusion.
III-A Overview
Our QMCIF framework is a patch-wise image fusion framework as shown in Fig. 1. Different from most existing works that perform color image fusion independently, our framework can handle all color channels simultaneously within a single fusion process. We first encode the input color images into their quaternion representations and then use the quaternion sparse decomposition model to jointly learn the optimal coefficient matrix and detail layer for each input image for high-precision focus detection. We employ the quaternion base-detail fusion strategy to fuse the color images in base and detail scales. In the detail-scale fusion process, the two-directional gradients of the detail layers are used to estimate the focus level of each input quaternion representation and select focused patches, generating a detail-scale fused result. In the base-scale fusion process, coefficient matrices are combined with the gradients of the corresponding detail layers to assess the focus level of each input quaternion representation and select focused patches to form the base-scale fused result. To further improve the fusion quality, we design a quaternion structural similarity refinement strategy to merge these fused results patch-by-patch to produce the final fused result using a quaternion weighted structural similarity index measure.
Figure 2: Our quaternion sparse decomposition (QSD) model.
III-B Quaternion sparse decomposition
III-B1 Model formulation
Given a partially focused color image $\bm{\mathrm{I}}$ $\in\mathbbm{R}^{M\times N\times 3}$ and its corresponding quaternion representation $\dot{\bm{\mathrm{I}}}$. Our goal is to decompose texture information from the static background as a strong focus measure. Our quaternion sparse decomposition (QSD) can be formulated as follows.
$\mathop{\arg\min}_{\dot{\bm{\mathrm{D}}},\dot{\bm{\mathrm{Z}}}}\ \ \sum_{k=1}^{K}{\|\dot{\bm{\mathrm{Z}}}_{k}\|}_{*}+\alpha\sum_{n=1}^{2}\|\nabla_{n}\dot{\bm{\mathrm{B}}}\|_{1}$ (3)
$+\beta\|\dot{\bm{\mathrm{D}}}\|_{1}+\lambda\|\dot{\bm{\mathrm{E}}}\|_{F}^{2}$
$s.t.\ \ \dot{\bm{\mathrm{I}}}=\dot{\bm{\mathrm{B}}}+\dot{\bm{\mathrm{D}}}+\dot{\bm{\mathrm{E}}},\ \mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}=\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}_{k}.$
In Eq. (3), ${\nabla}_{1}$ and ${\nabla}_{2}$ represent the vertical and horizontal gradients in the quaternion domain, respectively. The operator $\mathcal{R}(\cdot)$ extracts image patches of size $\sqrt{d}\times\sqrt{d}$ and stacks their vectorized forms into a quaternion matrix in $\mathbbm{H}^{d\times P}$. Column vectors can be divided into $K$ groups. A set of groups is denoted as $\{\mathcal{R}(\dot{\bm{\mathrm{B}}})_{1},\mathcal{R}(\dot{\bm{\mathrm{B}}})_{2},\cdots,\mathcal{R}(\dot{\bm{\mathrm{B}}})_{K}\}$, namely $\mathcal{R}(\dot{\bm{\mathrm{B}}})$ where $\mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}$ represents the $k^{th}$ group. $\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}}$ captures the low-rank property of $\mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}$. The nuclear norm of $\dot{\bm{\mathrm{Z}}}{{}_{k}}$ serves as a convex relaxation of the rank function. The positive parameters $\alpha$, $\beta$ and $\lambda$ balance the effects of each term in the optimization process.
As displayed in Fig. 2, our QSD converts a color image to its quaternion representation $\dot{\bm{\mathrm{I}}}$ that is then decomposed into a base-scale layer $\dot{\bm{\mathrm{B}}}$ with low-rank structural information, a detail-scale layer $\dot{\bm{\mathrm{D}}}$ with sparse details and a Gaussian noise layer $\dot{\bm{\mathrm{E}}}$. All layers are quaternion matrices belonging to $\mathbbm{H}^{M\times N}$. To fully utilize the focus information in $\dot{\bm{\mathrm{B}}}$, QSD uses nonlocal similarity to extract local patches in $\dot{\bm{\mathrm{B}}}$, stack these patches into column vectors, and partition them into several groups using the K-means clustering method. QSD learns a set of coefficient matrices $\dot{\bm{\mathrm{Z}}}_{k}$ from these groups using the quaternion dictionary $\dot{\bm{\mathrm{A}}}$ $\in\mathbbm{H}^{d\times L}$ containing $L$ atoms and concatenates these coefficient matrices to obtain $\dot{\bm{\mathrm{Z}}}$ $\in\mathbbm{H}^{L\times P}$. This ensures that patches within the same group are similar.
In general, the detail layer intuitively reflects the sharpness of edge pixels, as it exhibits a greater amplitude of pixel intensity in focused regions. However, in low-texture and boundary areas, the detail layer $\dot{\bm{\mathrm{D}}}$ often provides insufficient features. This may degrade fusion performance. To capture the focus information of smooth regions, coefficient matrix $\dot{\bm{\mathrm{Z}}}$ is introduced as a base-scale measure using quaternion sparse coding [21]. Obtaining the optimal $\dot{\bm{\mathrm{D}}}$ and $\dot{\bm{\mathrm{Z}}}$ simultaneously offers our QSD the greater flexibility than handcrafted filters.
III-B2 Optimization
To solve the optimization problem of our QSD in Eq. (3), we here introduce an iterative learning algorithm under the framework of the quaternion alternating direction method of multipliers [25, 29]. According to the definitions and lemmas in Section II, we write the Lagrangian function of our QSD in Eq. (3) as follows:
$\mathcal{L}$ $=\sum_{k=1}^{K}{\|\dot{\bm{\mathrm{J}}}_{k}\|}_{*}+\left\langle\dot{\bm{\mathrm{Y}}}{{}_{1,k}},\ \dot{\bm{\mathrm{Z}}}{{}_{k}}-\dot{\bm{\mathrm{J}}}{{}_{k}}\right\rangle+\frac{\mu}{2}\|\dot{\bm{\mathrm{Z}}}{{}_{k}}-\dot{\bm{\mathrm{J}}}{{}_{k}}\|_{F}^{2}$ (4)
$+\left\langle\dot{\bm{\mathrm{Y}}}{{}_{2,k}},\ \mathcal{R}(\dot{\bm{\mathrm{B}}}){{}_{k}}-\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}}\right\rangle+\frac{\mu}{2}\|\mathcal{R}(\dot{\bm{\mathrm{B}}}){{}_{k}}-\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}}\|_{F}^{2}$
$+\alpha\|\dot{\bm{\mathrm{G}}}_{1}\|_{1}+\left\langle\dot{\bm{\mathrm{Y}}}{{}_{3}},\ \dot{\bm{\mathrm{G}}}_{1}-\nabla_{1}\dot{\bm{\mathrm{B}}}\right\rangle+\frac{\mu}{2}\|\dot{\bm{\mathrm{G}}}_{1}-\nabla_{1}\dot{\bm{\mathrm{B}}}\|_{F}^{2}$
$+\alpha\|\dot{\bm{\mathrm{G}}}_{2}\|_{1}+\left\langle\dot{\bm{\mathrm{Y}}}{{}_{4}},\ \dot{\bm{\mathrm{G}}}_{2}-\nabla_{2}\dot{\bm{\mathrm{B}}}\right\rangle+\frac{\mu}{2}\|\dot{\bm{\mathrm{G}}}_{2}-\nabla_{2}\dot{\bm{\mathrm{B}}}\|_{F}^{2}$
$+\beta\|\dot{\bm{\mathrm{D}}}\|_{1}+\lambda\|\dot{\bm{\mathrm{E}}}\|_{F}^{2}+\left\langle\dot{\bm{\mathrm{Y}}}{{}_{5}},\ \dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}-\dot{\bm{\mathrm{D}}}-\dot{\bm{\mathrm{E}}}\right\rangle$
$+\frac{\mu}{2}\|\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}-\dot{\bm{\mathrm{D}}}-\dot{\bm{\mathrm{E}}}\|_{F}^{2}.$
[…]
Update ${\{\dot{\bm{\mathrm{J}}}{{}_{k}}\}}_{k=1}^{K}$. Fix ${\{\dot{\bm{\mathrm{Z}}}{{}_{k}}\}}_{k=1}^{K}$ and ${\{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}\}}_{k=1}^{K}$. Let $\dot{\bm{\mathrm{P}}}=$ $\dot{\bm{\mathrm{Z}}}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}}{\mu}$. $\dot{\bm{\mathrm{J}}}{{}_{k}}$ is minimized using the quaternion singular value thresholding method [26].
$\dot{\bm{\mathrm{J}}}{{}_{k}}$ $=\mathop{\arg\min}_{\dot{\bm{\mathrm{J}}}{{}_{k}}}\ \ {\|\dot{\bm{\mathrm{J}}}_{k}\|}_{*}+\frac{\mu}{2}\|\dot{\bm{\mathrm{J}}}{{}_{k}}-(\dot{\bm{\mathrm{Z}}}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}}{\mu})\|_{F}^{2}$ (5)
$=\mathop{\arg\min}_{\dot{\bm{\mathrm{J}}}{{}_{k}}}\ \ \frac{1}{\mu}{\|\dot{\bm{\mathrm{J}}}_{k}\|}_{*}+\frac{1}{2}\|\dot{\bm{\mathrm{J}}}{{}_{k}}-\dot{\bm{\mathrm{P}}}\|_{F}^{2}.$
Therefore, Eq. (5) can be solved using the Lemma 1.
Update ${\{\dot{\bm{\mathrm{Z}}}{{}_{k}}\}}_{k=1}^{K}$. Fix ${\{\dot{\bm{\mathrm{B}}}{{}_{k}}\}}_{k=1}^{K}$, ${\{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}\}}_{k=1}^{K}$ and ${\{\dot{\bm{\mathrm{Y}}}{{}_{2,k}}\}}_{k=1}^{K}$. Let $\dot{\bm{\mathrm{Q}}}=\dot{\bm{\mathrm{J}}}{{}_{k}}-\frac{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}}{\mu}$. The $\dot{\bm{\mathrm{Z}}}{{}_{k}}$-subproblem is reduced to:
$\mathop{\arg\min}_{\dot{\bm{\mathrm{Z}}}{{}_{k}}}\ \ \frac{\mu}{2}{\|\dot{\bm{\mathrm{Z}}}{{}_{k}}-\dot{\bm{\mathrm{Q}}}\|}_{F}^{2}+\frac{\mu}{2}\|\mathcal{R}(\dot{\bm{\mathrm{B}}}){{}_{k}}-\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{2,k}}}{\mu}\|_{F}^{2}.$ (6)
We set the gradient of Eq. (6) with respect to $\dot{\bm{\mathrm{Z}}}{{}_{k}}$ to zero. $\dot{\bm{\mathrm{Z}}}{{}_{k}}$ is explicitly updated as follows.
$\begin{split}\dot{\bm{\mathrm{Z}}}{{}_{k}}={(\dot{\bm{\mathrm{A}}}{{}^{\mathrm{H}}}\dot{\bm{\mathrm{A}}}+\dot{\bm{\mathrm{I}}}_{d})}^{-1}(\dot{\bm{\mathrm{A}}}{{}^{\mathrm{H}}}({\mathcal{R}(\dot{\bm{\mathrm{B}}})}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{2,k}}}{\mu})+\dot{\bm{\mathrm{Q}}}).\end{split}$ (7)
Update $\dot{\bm{\mathrm{G}}}{{}_{1}}$ and $\dot{\bm{\mathrm{G}}}{{}_{2}}$. Fix $\dot{\bm{\mathrm{B}}}$ and $\dot{\bm{\mathrm{Y}}}{{}_{3}}$. $\dot{\bm{\mathrm{G}}}{{}_{1}}$ is updated according to the subproblem:
[…]
The solution to Eq. (8) is obtained using the soft-thresholding method in Lemma 2. Similarly, $\dot{\bm{\mathrm{G}}}{{}_{2}}$ is updated in the same way as the $\dot{\bm{\mathrm{G}}}{{}_{1}}$-subproblem.
Update $\dot{\bm{\mathrm{B}}}$. To compute the solution of $\dot{\bm{\mathrm{B}}}$-subproblem, variables ${\{\dot{\bm{\mathrm{Z}}}{{}_{k}}\}}_{k=1}^{K}$, $\dot{\bm{\mathrm{G}}}_{1}$, $\dot{\bm{\mathrm{G}}}_{2}$, $\dot{\bm{\mathrm{D}}}$, $\dot{\bm{\mathrm{E}}}$, ${\{\dot{\bm{\mathrm{Y}}}{{}_{2,k}}\}}_{k=1}^{K}$, $\dot{\bm{\mathrm{Y}}}_{3}$, $\dot{\bm{\mathrm{Y}}}_{4}$, and $\dot{\bm{\mathrm{Y}}}_{5}$ are fixed. Let $\dot{\bm{\mathrm{M}}}{{}_{3,k}}=\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}}-\frac{\dot{\bm{\mathrm{Y}}}{{}_{2,k}}}{\mu}$. ${\{\dot{\bm{\mathrm{M}}}{{}_{3,k}}\}}_{k=1}^{K}$ is computed for each group. We reverse the extraction process of local quaternion image patches using the inverse operator ${\mathcal{R}}^{-1}$ applied to $\dot{\bm{\mathrm{M}}}{{}_{3}}$. It is the stacked version of ${\{\dot{\bm{\mathrm{M}}}{{}_{3,k}}\}}_{k=1}^{K}$. Let $\dot{\bm{\mathrm{M}}}_{4}=\dot{\bm{\mathrm{G}}}{{}_{1}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{3}}}{\mu}$, $\dot{\bm{\mathrm{M}}}_{5}=\dot{\bm{\mathrm{G}}}{{}_{2}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{4}}}{\mu}$, and $\dot{\bm{\mathrm{M}}}_{6}=\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{D}}}-\dot{\bm{\mathrm{E}}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{5}}}{\mu}$. The $\dot{\bm{\mathrm{B}}}$-subproblem can be rewritten as below.
$\dot{\bm{\mathrm{B}}}=$ $\mathop{\arg\min}_{\dot{\bm{\mathrm{B}}}}\ \ \frac{\mu}{2}\|\dot{\bm{\mathrm{B}}}-\mathcal{R}^{-1}({\dot{\bm{\mathrm{M}}}_{3}})\|_{F}^{2}+\frac{\mu}{2}\|\nabla_{1}\dot{\bm{\mathrm{B}}}-\dot{\bm{\mathrm{M}}}_{4}\|{{}_{F}^{2}}$ (9)
$+\frac{\mu}{2}\|\nabla_{2}\dot{\bm{\mathrm{B}}}-\dot{\bm{\mathrm{M}}}_{5}\|{{}_{F}^{2}}+\frac{\mu}{2}\|\dot{\bm{\mathrm{B}}}-\dot{\bm{\mathrm{M}}}_{6}\|{{}_{F}^{2}}.$
We set the derivatives of Eq. (9) with respect to $\dot{\bm{\mathrm{B}}}$ to zero. $\dot{\bm{\mathrm{B}}}$ is updated using quaternion fast Fourier transform $\mathcal{F}(\cdot)$ under the quaternion periodic boundary condition[29]. Let $\dot{\bm{\mathrm{\Sigma}}}$ denote the result of $\mathcal{F}({\mathcal{R}}^{-1}({\dot{\bm{\mathrm{M}}}_{3}}))+\mathcal{F}(\nabla_{1}^{\mathrm{T}})\cdot\mathcal{F}(\dot{\bm{\mathrm{M}}}_{4})+\mathcal{F}(\nabla_{2}^{\mathrm{T}})\cdot\mathcal{F}(\dot{\bm{\mathrm{M}}}_{5})+\mathcal{F}(\dot{\bm{\mathrm{M}}}_{6})$. $\dot{\bm{\mathrm{B}}}$ is updated as below.
$\begin{split}\dot{\bm{\mathrm{B}}}=\mathcal{F}{{}^{-1}}({\frac{\dot{\bm{\mathrm{\Sigma}}}}{\mathcal{F}(\nabla_{1}^{\mathrm{T}}\nabla_{1}+\nabla_{2}^{\mathrm{T}}\nabla_{2})+2}}).\end{split}$ (10)
Update $\dot{\bm{\mathrm{D}}}$. We fix $\dot{\bm{\mathrm{B}}}$, $\dot{\bm{\mathrm{E}}}$, and $\dot{\bm{\mathrm{Y}}}_{4}$ to minimize $\dot{\bm{\mathrm{D}}}$. The $\dot{\bm{\mathrm{D}}}$-subproblem can be rewritten as follows:
$\begin{split}\dot{\bm{\mathrm{D}}}=\mathop{\arg\min}_{\dot{\bm{\mathrm{D}}}}\ \ \frac{\beta}{\mu}{\|\dot{\bm{\mathrm{D}}}\|}_{1}+\frac{1}{2}\|\dot{\bm{\mathrm{D}}}-\dot{\bm{\mathrm{M}}}_{7}\|_{F}^{2}.\end{split}$ (11)
Eq. (11) can also be solved using Lemma 2.
Update $\dot{\bm{\mathrm{E}}}$. Fix $\dot{\bm{\mathrm{B}}}$, $\dot{\bm{\mathrm{D}}}$, and $\dot{\bm{\mathrm{Y}}}_{5}$. Let $\dot{\bm{\mathrm{M}}}_{8}=\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{D}}}-\dot{\bm{\mathrm{B}}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{5}}}{\mu}$. The solution of the $E$-subproblem is obtained by setting the gradient values of $\dot{\bm{\mathrm{E}}}$’s subproblem with respect to $\dot{\bm{\mathrm{E}}}$ to zero. $\dot{\bm{\mathrm{E}}}$ is updated as below.
$\dot{\bm{\mathrm{E}}}={(2\lambda+\mu\dot{\bm{\mathrm{I}}}_{d})}^{-1}(\mu\dot{\bm{\mathrm{M}}}_{8}),$ (12)
The optimization process of the QSD model is shown in Algorithm 1.
Input: The quaternion representation of the source image $\dot{\bm{\mathrm{I}}}$, the parameters $\mu$, $\lambda$, $\alpha$ and $\beta$, the iteration number $t$.
Output: Optimal quaternion coefficient matrix $\dot{\bm{\mathrm{Z}}}$ and detail layer $\dot{\bm{\mathrm{D}}}$
1 Initialize ${\{\dot{\bm{\mathrm{J}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$,${\{\dot{\bm{\mathrm{Z}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$, ${\{\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{0}}\}}{{}_{k=1}^{K}}$,${\{\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{0}}\}}{{}_{k=1}^{K}}$, $\dot{\bm{\mathrm{G}}}{{}_{1}^{0}}$, $\dot{\bm{\mathrm{G}}}{{}_{2}^{0}}$, $\dot{\bm{\mathrm{B}}}{{}^{0}}$, $\dot{\bm{\mathrm{D}}}{{}^{0}}$, $\dot{\bm{\mathrm{E}}}{{}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{3}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{4}^{0}}$ and $\dot{\bm{\mathrm{Y}}}{{}_{5}^{0}}$;
2while not converged do
3       for $k=1,\cdots,K$ do
4             Fix other variables and compute $\dot{\bm{\mathrm{J}}}{{}_{k}^{t}}$ using Lemma 1;
5            Fix other variables and compute $\dot{\bm{\mathrm{Z}}}{{}_{k}^{t}}$ using Eq. (7);
6       end for
7      Fix other variables and solve $\dot{\bm{\mathrm{G}}}{{}_{1}^{t}}$ and $\dot{\bm{\mathrm{G}}}{{}_{2}^{t}}$ using Lemma 2 respectively;
8      Fix other variables and solve $\dot{\bm{\mathrm{B}}}{{}^{t}}$ using Eq. (10);
9      Fix other variables and solve $\dot{\bm{\mathrm{D}}}{{}^{t}}$ using Lemma 2;
10      Fix other variables and solve $\dot{\bm{\mathrm{E}}}{{}^{t}}$ using Eq. (12);
11      for $k=1,\cdots,K$ do
12             $\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{t}}=\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{t-1}}+\mu(\dot{\bm{\mathrm{Z}}}{{}_{k}^{t}}-\dot{\bm{\mathrm{J}}}{{}_{k}^{t}})$;
13            $\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{t}}=\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{t-1}}+\mu(\mathcal{R}(\dot{\bm{\mathrm{B}}}){{}_{k}^{t}}-\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}^{t}})$;
14       end for
15
16      ${\dot{\bm{\mathrm{Y}}}{{}_{3}^{t}}}={\dot{\bm{\mathrm{Y}}}{{}_{3}^{t-1}}}+\mu(\dot{\bm{\mathrm{G}}}{{}_{1}^{t}}-\nabla_{1}\dot{\bm{\mathrm{B}}}{{}^{t}})$;
17      ${\dot{\bm{\mathrm{Y}}}{{}_{4}^{t}}}={\dot{\bm{\mathrm{Y}}}{{}_{4}^{t-1}}}+\mu(\dot{\bm{\mathrm{G}}}{{}_{2}^{t}}-\nabla_{2}\dot{\bm{\mathrm{B}}}{{}^{t}})$;
18      ${\dot{\bm{\mathrm{Y}}}{{}_{5}^{t}}}={\dot{\bm{\mathrm{Y}}}{{}_{5}^{t-1}}}+\mu(\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}{{}^{t}}-\dot{\bm{\mathrm{D}}}{{}^{t}}-\dot{\bm{\mathrm{E}}}{{}^{t}})$;
19      $\mu=\min{\{{10}^{6},\mu*1.1\}}$;
20      $t$longleftarrow $t+1$;
21 end while
Algorithm 1 Quaternion sparse decomposition
III-C Quaternion base-detail fusion strategy
This subsection proposes a quaternion base-detail fusion (QBDF) strategy to achieve high-quality color image fusion in base and detail scales. Fig. 3 illustrates our QBDF strategy. Given quaternion representations ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{2}$ $\in\mathbbm{H}^{2\times M\times N}$ of two input color images, our QBDF strategy first uses the QSD model to learn their detail-scale layers ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{2}$ and coefficient matrices ${\{\dot{\bm{\mathrm{Z}}}^{(i)}\}}_{i=1}^{2}$, and then performs two-scale image fusion.
To fully utilize the complementary features of coefficient matrices and detail-scale layers, QBDF strategy introduces a two-scale fusion process to obtain base-scale and detail-scale fusion results at the patch level. The patch-wise fusion rule is adopted to reduce the sensitivity to the undetermined pixels directly generated by focus detection. First, we define two-scale focus measures as below.
$l_{b,j}=\mathop{\arg\max}_{i}\ \ \sum_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}_{1}+\theta{\|\dot{\bm{\mathrm{Z}}}_{j}^{(i)}\|}_{2},$ (13)
$l_{d,j}=\mathop{\arg\max}_{i}\ \ \sum_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}_{1},$ (14)
where $l_{b}$ and $l_{d}$ denote the patch-wise focus maps of base-scale and detail-scale fused results respectively, $j$ defines the center spatial location of each patch and parameter $\theta$ is here set to 1 by default. $\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}$ represents the $j^{th}$ patch of $\dot{\bm{\mathrm{D}}}^{(i)}$ and $\dot{\bm{\mathrm{Z}}}{{}^{(i)}_{j}}$ corresponds to the $j^{th}$ column vector of $\dot{\bm{\mathrm{Z}}}^{(i)}$. The detail-scale focus measure uses the two-directional gradients of the detail-scale layers. The gradients that intuitively represent variations in the quaternion pixels can substantially reflect the focus levels of color images. Based on the greater amplitude of Eq. (14), $l_{d,j}$ is assigned with the corresponding $i$ in $j^{th}$ patch.
Figure 3: Flowchart of our quaternion base-detail fusion (QBDF) and quaternion structural similarity refinement (QSSR) strategies.
Although the detail-scale focus map $l_{d}$ effectively corrects the inaccurate focus map using the neighboring information for boundary and small focused regions, it may produce visual artifacts and fail to handle low-gradient background areas. This limitation can lead to significant inconsistencies in the detail-scale fused result. To address this issue, the base-scale focus measure integrates the gradients of the detail layers with their corresponding coefficient matrices. This ensures local consistency and mitigates spatial artifacts caused by the detail-scale focus map. Based on the greater amplitude of Eq. (13), $l_{b,j}$ is assigned with the corresponding $i$ in $j^{th}$ patch.
The focused patches are obtained and aggregated into detail-scale and the base-scale fused results ${\dot{\bm{\mathrm{P}}}}^{(1)}$ and ${\dot{\bm{\mathrm{P}}}}^{(2)}$ using the following two-scale patch-wise fusion rule:
$\dot{\bm{\mathrm{P}}}{{}^{(1)}_{j}}=\left\{\begin{aligned} &\dot{\bm{\mathrm{I}}}{{}^{(1)}_{j}},&&l_{d,j}=1\\ &\dot{\bm{\mathrm{I}}}{{}^{(2)}_{j}},&&l_{d,j}=2\end{aligned}\right.$ (15)
$\dot{\bm{\mathrm{P}}}{{}^{(2)}_{j}}=\left\{\begin{aligned} &\dot{\bm{\mathrm{I}}}{{}^{(1)}_{j}},&&l_{b,j}=1\\ &\dot{\bm{\mathrm{I}}}{{}^{(2)}_{j}},&&l_{b,j}=2\end{aligned}\right.$ (16)
III-D Quaternion structural similarity refinement strategy
To obtain a high-quality fused result, we introduce a quaternion structural similarity refinement (QSSR) strategy to refine the base-scale and detail-scale fused patches of ${\{\dot{\bm{\mathrm{P}}}^{(i)}\}}_{i=1}^{2}$ using a selection rule with a maximum quaternion image structural similarity value. We assume that the patch to be selected should be most similar to the corresponding region of the source image with the largest focus information. Therefore, we propose a quaternion weighted structural similarity (QWSSIM) strategy to refine the distorted patches that are wrongly detected as focused ones based on the theory of Definition 5. QWSSIM is defined in the following equation.
$QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)})=\sum_{i=1}^{2}{\tau}_{i,j}QSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)},{\dot{\bm{\mathrm{I}}}}_{j}^{(i)}),$ (17)
$QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)})=\sum_{i=1}^{2}{\tau}_{i,j}QSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)},{\dot{\bm{\mathrm{I}}}}_{j}^{(i)}),$
where $j\in[1,P]$ denotes the center spatial location of each patch where $P$ denotes the number of overlapping patches. ${\tau}_{i,j}$ is an assigned weight to $j^{th}$ quaternion patch of $i^{th}$ input quaternion representation. The patches selected from the two-scale fused results by our QBDF strategy should be quite similar to the corresponding focused patches of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}^{2}_{i=1}$. In other words, each local quaternion patch of the final fused result has a maximum QWSSIM value with the corresponding region of ${\{\dot{\bm{\mathrm{I}}}\}}^{2}_{i=1}$. ${\tau}_{i,j}$ is significantly larger when the corresponding region of ${\dot{\bm{\mathrm{I}}}}^{(i)}$ is focused. Fig. 3 illustrates the flowchart of our QSSR strategy. Our QSSR strategy further compares two-scale fused results with the input quaternion representations patch-by-patch to assess the patch-wise structural similarity using our QWSSIM measure. Patches with higher QWSSIM values are selected to form the final fused result.
To adaptively compute ${\tau}_{i,j}$, we utilize the detail-scale measure to generate each weight map with the size of $n\times P$ where n depends on the number of inputs. The detail-scale measure is enhanced using a Laplacian function $\phi(x)=1-e^{-\frac{x}{\gamma}}$ where $\gamma$ is $=0.2$ by default. Then the weight map is formulated in the following equation.
${\tau}_{i,j}=\frac{\phi({({\sum}_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}{{}_{1}})})}{\sum_{i}\phi(({\sum}_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}{{}_{1}}))}.$ (18)
We apply a maximum selection rule based on the QWSSIM measure at the patch level. It effectively corrects defocused regions in the two-scale fused images. Additionally, this strategy eliminates visual artifacts caused by inaccurate focus maps. The selection rule is defined as follows.
${\bm{\mathrm{W}}}_{j}=\left\{\begin{aligned} &1,&QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)})>QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)})\\ &0,&QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)})\leq QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)}).\end{aligned}\right.$ (19)
The weight matrix $\bm{\mathrm{W}}$ denotes a refined patch-wise focus map. When ${\bm{\mathrm{W}}}_{j}=1$, we select the $j^{th}$ patch from ${\dot{\bm{\mathrm{P}}}}^{(1)}$. Otherwise, we choose the $j^{th}$ patch from ${\dot{\bm{\mathrm{P}}}}^{(2)}$. The final fused result is obtained at the patch level as below:
$\dot{\bm{\mathrm{F}}}_{j}=\left\{\begin{aligned} &{\dot{\bm{\mathrm{P}}}}_{j}^{(1)},&{\bm{\mathrm{W}}}_{j}=1\\ &{\dot{\bm{\mathrm{P}}}}_{j}^{(2)},&{\bm{\mathrm{W}}}_{j}=0.\end{aligned}\right.$ (20)
The patches are aggregated to form a quaternion fused result ${\dot{\bm{\mathrm{F}}}}$. After the inverse transformation from the quaternion domain into the real domain, we obtain a fused color image $\bm{\mathrm{F}}$ $\in\mathbbm{R}^{M\times N\times 3}$.
III-E Extension to multiple color image fusion
Fig. 1 displays our quaternion multi-focus color image fusion framework to fuse two partially focused color images. Our framework is flexible to be extended into fusing multiple color images. Let ${\{{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ denote $n$ input color images ($n\geq 2$). The multi-focus color image fusion performed on ${\{{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ is decribed in the following process.
Focus detection. We first encode multiple color images into their quaternion representations ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$. The QSD model is then used to decompose them into quaternion coefficient matrices ${\{\dot{\bm{\mathrm{Z}}}^{(i)}\}}_{i=1}^{n}$ and quaternion detail layers ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ in the iterative learning process.
QBDF strategy. Our QBDF strategy uses two-scale focus measures to obtain the fused results. The detail-scale focus measure uses two-directional gradients of $\dot{\bm{\mathrm{D}}}^{(i)}$ while the base-scale focus measure integrates the gradients of $\dot{\bm{\mathrm{D}}}^{(i)}$ with their corresponding $\dot{\bm{\mathrm{Z}}}^{(i)}$ for each quaternion representation. The base-scale and detail-scale measures are easily extended to ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ ($n\geq 2$) using Eqs. (13) and (14). We select focused patches to obtain base-scale and detail-scale fused results ${\{\dot{\bm{\mathrm{P}}}^{(i)}\}}_{i=1}^{2}$ based on Eqs. (15) and (16) similarly.
QSSR strategy. Our QSSR strategy compares each patch from base-scale and detail-scale fused results with input quaternion representations respectively using the QWSSIM. The patches with a higher QWSSIM value are selected as focused ones to form a final fused result. Therefore, the QBDF and QSSR strategies are easily extended to multiple color images. The detailed process is presented in Algorithm 2.
Input: Input $n$ color images ${\{\bm{\mathrm{I}}^{(i)}\}}{{}_{i=1}^{n}}$ ($n\geq 2$)
Output: The fused image $\bm{\mathrm{F}}$
1 for $i=1,\cdots,n$ do
2       Transform $\bm{\mathrm{I}}^{(i)}\in\mathbbm{R}$ into $\dot{\bm{\mathrm{I}}}^{(i)}$ in the quaternion domain;
3      Solve the quaternion sparse coefficient matrix ${\dot{\bm{\mathrm{Z}}}}^{(i)}$ and the quaternion detail layer ${\dot{\bm{\mathrm{D}}}}^{(i)}$ using Algorithm 1;
4 end for
5Divide each quaternion representation of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ and ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ into $P_{1}$ overlapping patches respectively;
6Fuse the detail-scale result ${\dot{\bm{\mathrm{P}}}}^{(1)}$ using Eq. (14) ;
7Divide each quaternion representation of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ and ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ into $P_{2}$ overlapping patches respectively;
8Fuse the base-scale result ${\dot{\bm{\mathrm{P}}}}^{(2)}$ using Eq. (13);
9Divide ${\dot{\bm{\mathrm{P}}}}^{(1)}$ and ${\dot{\bm{\mathrm{P}}}}^{(2)}$ into $P$ overlapping patches respectively;
10Divide each quaternion representation of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ and ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ into $P$ overlapping patches respectively;
11Compute $QWSSIM$ values at the patch level using Eq. (17);
12Obtain the refined focus map and fused patches at the patch level using Eqs. (19) and (20);
Aggregate the patches of ${\{\dot{\bm{\mathrm{F}}}_{j}\}}_{j=1}^{P}$ to obtain the final result $\bm{\mathrm{F}}$;
Algorithm 2 Quaternion Multi-focus Color Image Fusion
IV Experiments
To evaluate the fusion performance of our framework, we conduct experiments on three public datasets. Section A presents experimental settings like competing methods, datasets, and evaluation metrics. Sections B and C discuss the convergence and optimal parameters of the QSD model. To demonstrate the superiority of our method, Sections D and E present experimental results compared with other state-of-the-art approaches. Section F presents ablation studies on the mffw dataset [30] to verify the effectiveness of our QSD model, QBDF and QSSR strategies.
IV-A Experimental settings
In our experiments, the patch size is set as $\sqrt{d}\times\sqrt{d}$ where $\sqrt{d}=8$ by default in the QSD model and base-scale fusion process. In the detail-scale fusion and the QSSR process, $\sqrt{d}$ is set to $5\times 10^{-3}MN$ [18], where $M$ and $N$ denote the height and width of the input image. The stopping criterion of QSD model is $51e-$ 5 in Algorithm 1. All experiments are performed using Matlab 2016 on a workstation with a 2.90GHz Intel Core CPU and 16GB memory.
Comparison methods. Nine representative multi-focus image fusion algorithms are selected as the comparison methods. They include sparse representation-based methods (SR) [10] as the baseline method, Guide-filter-based method (GFDF) [11], Multi-scale gradients-based method (MGIMF) [16], GAN-based method (MFIF-GAN) [31], Higher-order SVD-based method (QHOSVD) [19], unsupervised zero-shot based method (ZMFF) [32], Small-area-aware method (SAMF) [4], End-to-End based network (DBMFIF) [7], General-Image-Fusion-based (TCMOA) [33]. SR belongs to channel-wise processing-based MCIF methods. GFDF, MGIMF, and SAMF are grayscale conversion-based MCIF methods. QHOSVD is a quaternion domain-based MCIF method. MFIF-GAN, ZMFF, DBMFIF, and TCMOA are deep-learning-based MCIF methods.
Datasets. In our experiments, we utilize three public datasets without ground truth: the lytro dataset [34], mffw dataset [30], and MFI-WHU dataset [35]. The lytro dataset contains 20 pairs of samples for two-color image fusion and 4 pairs of samples for multiple-color image fusion. The mffw dataset includes 13 pairs of samples for two-color image fusion and 6 pairs of image data for multiple-source fusion. The MFI-WHU dataset provides 30 pairs of samples for two-color image fusion.
The images in the lytro dataset [34] are well-registered and exhibit high levels of details. It aims to evaluate focus information preservation in ideal conditions. The images in the mffw dataset [30] are affected by severe defocus spread effects with large misalignment and contain both low-detail and high-detail regions. The performance on the mffw dataset provides a robust evaluation of focus information preservation and artifact removal in complex real-world scenarios. The MFI-WHU dataset [35] includes input images with tiny blurred objects. These pose challenges for small-area focus detection and fusion.
Metrics. Following [2], six objective metrics are adopted to evaluate the performance of two-color image fusion quantitatively. They are the normalized mutual information $Q_{MI}$, gradient-based metric $Q_{G}$, phase congruency-based metric $Q_{P}$, structural similarity-based metric $Q_{E}$, $Q_{Y}$ and human perception-based metric $Q_{CB}$. $Q_{MI}$ quantifies the intensity-based mutual information transferred from source images to the fused output. $Q_{G}$ and $Q_{P}$ assess the preservation of gradient information and salient feature details. $Q_{Y}$ evaluates the structural preservation of the fused image in terms of luminance, contrast, and texture patterns while $Q_{E}$ extends $Q_{Y}$ by considering edge consistency in the fused result. $Q_{CB}$ measures the preservation of contrast variations from source images in the fused image.
IV-B Convergence Analysis
The convergence analysis of the QSD model is provided in this subsection. Fig. 4 displays the evolution curves of the relative difference versus iterations of QSD model on the samples of lytro [34], mffw [30] and MFI-WHU datasets [35]. The relative difference is calculated by the maximum value between $\|\dot{\bm{\mathrm{Z}}}{{}^{t+1}}-\dot{\bm{\mathrm{Z}}}{{}^{t}}\|{{}_{\infty}}$ and $\|\dot{\bm{\mathrm{D}}}{{}^{t+1}}-\dot{\bm{\mathrm{D}}}{{}^{t}}\|{{}_{\infty}}$. $\dot{\bm{\mathrm{Z}}}{{}^{t+1}}$ and $\dot{\bm{\mathrm{Z}}}{{}^{t}}$ denote the coefficient matrices at successive $t^{th}$ and ${(t+1)}^{th}$ iterations respectively. $\dot{\bm{\mathrm{D}}}{{}^{t+1}}$ and $\dot{\bm{\mathrm{D}}}{{}^{t}}$ denote the detail layers at successive $t^{th}$ and ${(t+1)}^{th}$ iterations respectively. The $\infty$ norm means the largest magnitude among each element of a vector. To improve the QSD speed, $\dot{\bm{\mathrm{B}}}{{}^{0}}$ is initialized as $\frac{\mathcal{F}(\dot{\bm{\mathrm{I}}})}{\mathcal{F}(\nabla_{1}^{T}\nabla_{1}+\nabla_{2}^{T}\nabla_{2})+1}$ and $\dot{\bm{\mathrm{D}}}{{}^{0}}$ is set to $\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}{{}^{0}}$ As displayed in Fig. 4, QSD is convergent and its total number of iterations is 20. This can sufficiently guarantee convergence.
Figure 4: Convergence analysis of QSD model on lytro, mffw and MFI-WHU datasets.
IV-C Parameter Discussions
Here, we discuss how three regularization parameters $\alpha$, $\beta$, and $\lambda$ affect the fusion performance of the QSD model. First, we show the visual effects of various values of $\alpha$, $\beta$ and $\lambda$. Then we comprehensively analyze and determine parameter settings.
Figure 5: The fusion performance with different parameters on the lytro and the mffw datasets. (a) $Q_{E}+Q_{CB}$ results when parameters $\alpha$ and $\beta$ vary on the mffw dataset. (b) The $Q_{MI}$ value with respect to parameter $\lambda$ on both lytro and mffw datasets.
In our QSD model, parameter $\alpha$ controls the base-scale layer while parameter $\beta$ handles the detail-scale layer. The final fused image not only preserves the most detailed parts of the inputs but also maintains the consistency of the background.
To handle the three regularization parameters in our QSD model, we fix parameter $\lambda=0.05$ and adjust the others each time. $\alpha$ and $\beta$ were sequentially set values of [0.5, 1, 1.5, 2, 2.5, 3]. Then $\alpha$ and $\beta$ were set to 1.5 and 2 and the variation values of $\lambda$ were sequentially set values of [0.05, 0.06, 0.07, 0.08, 0.09, 0.1]. To effectively estimate the fusion performance, the combination index $Q_{E}+Q_{CB}$ approximates the fused images in terms of detail and base parts. Since $\lambda$ is not remarkably sensitive to the final fusion quality, index $Q_{MI}$ is used to quantitatively analyze the fused results.
To quantitatively analyze the sensitivity of QSD versus these parameters, Fig. 5 displays the $Q_{E}+Q_{CB}$ results of QSD with the samples of mffw dataset. In Fig. 5 (a), $\alpha$ and $\beta$ are set to 1.5 and 2 on mffw dataset and $\beta$ are set to 0.5 on lytro and MFI-WHU datasets. $\lambda=5e-2$ is set for all datasets in Fig. 5 (b).
TABLE II: Performance comparison of multi-focus image fusion methods in two-color image fusion.
Dataset Methods $Q_{MI}\uparrow$ $Q_{G}\uparrow$ $Q_{P}\uparrow$ $Q_{E}\uparrow$ $Q_{Y}\uparrow$ $Q_{CB}\uparrow$
lytro [34] SR [10] 1.0642 0.7445 0.8168 0.8797 0.9688 0.7756
GFDF [11] 1.1194 0.7524 0.8335 0.8799 0.9826 0.8021
MGIMF [16] 1.1154 0.7428 0.8106 0.8419 0.9857 0.7911
QHOSVD [19] 1.0540 0.7491 0.8310 0.8800 0.9769 0.7839
SAMF [4] 1.1781 0.7602 0.8412 0.8797 0.9877 0.8014
MFIF-GAN [31] 1.0945 0.7179 0.8307 0.8787 0.9770 0.7976
ZMFF [32] 0.8796 0.7013 0.7830 0.8678 0.9313 0.7399
DBMFIF [7] 1.0589 0.7480 0.8398 0.8802 0.9651 0.7775
TCMOA [33] 0.9926 0.7401 0.8188 0.8782 0.9574 0.7597
Ours 1.1656 0.7613 0.8486 0.8813 0.9896 0.8056
mffw [30] SR [10] 0.9181 0.6278 0.5500 0.8078 0.7844 0.6503
GFDF [11] 1.0878 0.7323 0.7630 0.8242 0.9522 0.7426
MGIMF [16] 1.0529 0.7309 0.7363 0.8096 0.9507 0.7338
QHOSVD [19] 0.7224 0.7249 0.7383 0.8187 0.8221 0.6050
SAMF [4] 1.0863 0.7310 0.6995 0.7974 0.9332 0.7101
MFIF-GAN [31] 1.0681 0.7320 0.7550 0.8239 0.9735 0.7558
ZMFF [32] 0.7728 0.6651 0.6476 0.7985 0.8775 0.6770
DBMFIF [7] 0.8703 0.6959 0.6620 0.8169 0.9048 0.6642
TCMOA [33] 0.7545 0.6034 0.5307 0.7874 0.7681 0.6475
Ours 1.0887 0.7338 0.7579 0.8275 0.9569 0.7450
MFI-WHU [35] SR [10] 1.1329 0.7294 0.7952 0.8424 0.9850 0.8236
GFDF [11] 1.1565 0.7337 0.8011 0.8416 0.9891 0.8312
MGIMF [16] 1.1269 0.7151 0.7782 0.7973 0.9825 0.8157
QHOSVD [19] 1.1507 0.7299 0.8008 0.8397 0.9869 0.8291
SAMF [4] 1.2155 0.7310 0.7980 0.8366 0.9890 0.8268
MFIF-GAN [31] 1.1648 0.7343 0.7994 0.8410 0.9890 0.8304
ZMFF [32] 0.6933 0.6342 0.6705 0.7915 0.8540 0.6860
DBMFIF [7] 1.0461 0.7219 0.7933 0.8424 0.9523 0.7900
TCMOA [33] 0.8635 0.6389 0.7026 0.7748 0.9139 0.7655
Ours 1.1743 0.7361 0.8017 0.8433 0.9887 0.8318
Figure 6: Two partially focused color image fusion on $lytro$, $mffw$ and $MFI-WHU$ datasets.
IV-D Two color image fusion
This subsection discusses the quantitative and qualitative results to assess fusion quality of two-color image fusion.
Quantitative evaluation. Table II provides the average performance of various competing algorithms on three public datasets. Six evaluation metrics are employed to assess the quality of the fused images, with higher values indicating better performance. In this paper, the highest evaluation scores are highlighted in bold.
On the lytro dataset [34], our framework achieves the highest metric values compared to the other competing methods across gradient-based, structural similarity-based, and human-perception-based evaluation metrics. This improvement is attributed to accurate focus detection capability of our method. However, our framework is slightly worse than SAMF on the metric of $Q_{MI}$. This is because our patch-wise fusion strategy tends to preserve less image intensity information during the quaternion fusion and refinement process compared to the pixel-wise fusion strategy.
Since the images on the mffw dataset are affected by severe defocus spread effects and contain more challenging low-detail regions [30], all competing methods perform worse on this dataset compared to lytro and MFI-WHU datasets. Our method achieves the highest fusion performance on $Q_{G}$, and $Q_{E}$ metrics on the mffw dataset, highlighting its superior capability in focus detection and detail preservation. In the $Q_{MI}$ results, our framework slightly outperforms GFDF thanks to the effective artifact removal of our QSSR strategy. On the metrics of $Q_{Y}$ and $Q_{CB}$, our framework is worse only than MFIF-GAN. This demonstrates the advantages of fusion and refinement provided by the GAN-based MCIF algorithm in mitigating defocus spread effects [31].
On the MFI-WHU dataset, our method achieves the best fusion performance in $Q_{G}$ and $Q_{E}$ metrics, underscoring its effectiveness in detecting focus regions from small areas, particularly when compared to SAMF. However, in the $Q_{Y}$ metric, our method performs slightly worse than GFDF and MFIF-GAN since both methods exhibit strong focus detection capabilities.
Extensive experiments on various datasets demonstrate our framework significantly outperforms QHOSVD in terms of detail preservation and artifact removal under real-world challenging scenarios.
Visual evaluation. To evaluate the effectiveness of detail preservation and artifact removal, our method is compared with nine state-of-the-art methods. The red and yellow bounding boxes highlight areas with quality degradation. They are enlarged to better illustrate differences in the focused and defocused regions across the input images and output fused results.
The first two rows of Fig. 6 present the fused results of all competing methods on the lytro dataset. The red bounding boxes reveal that the results of SR, MGIMF, SAMF, and QHOSVD exhibit significant blurring along the edges. MGIMF shows spatial artifacts in the regions within the yellow bounding box. This is because the suboptimal decision maps are generated by multi-scale hand-designed filters. Our framework generates a pleasant fused result with clear objects and without any artifacts.
The third and the fourth rows of Fig. 6 illustrate the fused results of all competing methods on the mffw dataset. The red bounding boxes indicate that the outputs of SR, GFDF, and QHOSVD contain numerous blurred pixels. Similarly, the yellow bounding box highlights pseudo-edges and ghosts in the boundary regions of the MGIMF result. Furthermore, the outputs of SR, GFDF, MFIF-GAN, MGIMF, and QHOSVD suffer from varying degrees of color distortion and spatial artifacts in low-gradient regions. Our framework gives the most satisfactory fused result because our framework sufficiently handles the low-grafient regions and prevents visual artifacts for uncertain boundary areas.
The last two rows of Fig. 6 show the fused images produced by all comparative methods on the MFI-WHU dataset. The input images contain small and near blurred areas. This makes it difficult to estimate the focus levels and accurately identify focused regions. As can be seen in the red and yellow bounding boxes of the results, it is evident that SR, GFDF, MFIF-GAN, MGIMF, QHOSVD, SAMF and TCMOA produce unclear regions such as electric wires and poles. In contrast, our method demonstrates superior accuracy in detecting and preserving these focused regions.
IV-E Multiple color image fusion
To verify the superiority of our method in multiple-color image fusion, we conduct the experiment on the mffw3 dataset.
Figure 7: Three partially focused color image fusion on $mffw3$ dataset. The source 1 image is a near-focus image. The source 2 data is a middle-focused image. The source 3 data is a far-focused image.
Qualitative evaluation. Since TCMOA [33] does not provide a pre-trained model, only nine methods are included in this comparison. The deep learning-based methods fuse the multiple color images according to [36].
As displayed in Fig. 7, the experimental results show that these comparison methods perform poorly in the multiple-color image fusion process. GFDF, MGIMF, ZMFF and SAMF contain noticeable visual artifacts in the background areas. In the red bounding boxes of the QHOSVD and MGIMF results, significant blurring is observed, particularly in the areas of the wings. Similarly, in the yellow bounding box of the SAMF result, the fused image is obstructed by defocused pixels within the patterns. SR shows obvious artifacts and color distortion. This may be because SR fails to provide accurate focus level for focus detection and it is highly sensitive to large misregistration and color differences in the source images. Fig. 7 indicates the superiority of our framework in terms of focus detection and artifact removal in foreground, background and boundary regions.
IV-F Ablation study
Effectiveness of QSD. To verify the advantages of the QSD model, we perform the ablation experiments on the mffw dataset (the first 10 pairs of samples) with two input color images. Different settings of parameters in Eq. (3) are provided, we investigate how the different terms of QSD influence the overall fusion performance. Specifically, three different settings of $\alpha=0$, $\beta=0$, $\lambda=0$ and $\alpha\neq 0$, $\beta\neq 0$, $\lambda\neq 0$ are particularly considered in this experiment. Table III gives the quantitative fusion results with different settings of QSD. As can be observed, the different settings of terms can influence the fusion performance at different degrees. The QSD model provides the best quantitative results. Therefore, this experiment further verifies the effectiveness of the $\alpha$ term, $\beta$ term and $\lambda$ term in Eq. (3).
TABLE III: Quantitative evaluation on the different settings of the QSD model, the QBDF and the QSSR strategies.
Settings $Q_{MI}\uparrow$ $Q_{G}\uparrow$ $Q_{P}\uparrow$ $Q_{E}\uparrow$ $Q_{CB}\uparrow$
Parameter $\bm{\alpha=0}$ 1.0580 0.7359 0.7488 0.8240 0.7298
$\bm{\beta=0}$ 1.0052 0.7338 0.7282 0.8279 0.6980
$\bm{\lambda=0}$ 1.0140 0.7324 0.7432 0.8259 0.7127
QSD 1.0685 0.7374 0.7832 0.8234 0.7508
Patch size $\bm{5\times 5}$ 1.0503 0.7376 0.7277 0.8263 0.7115
$\bm{10\times 10}$ 1.0401 0.7367 0.7567 0.8276 0.7276
$\bm{8\times 8}$ 1.0685 0.7374 0.7832 0.8234 0.7508
Fusion QBDF-Base-scale 0.9312 0.7021 0.7113 0.8099 0.6830
QBDF-Detail-scale 1.0388 0.7357 0.7457 0.8255 0.7228
QSSR 1.0685 0.7374 0.7832 0.8234 0.7508
Weight $\tau$ Same 1.0690 0.7346 0.7556 0.8183 0.7340
Adaptive 1.0685 0.7374 0.7832 0.8234 0.7508
Representation Real 0.9555 0.7082 0.6849 0.8043 0.6748
Quaternion 1.0685 0.7374 0.7832 0.8234 0.7508
Table III gives the quantitative fusion results with different patch sizes of the QSD model. As can be seen, the size $8\times 8$ provides the best quantitative results among five metrics.
Effectiveness of QBDF and QSSR. To verify the advantages of the QBDF and the QSSR strategies, we perform the ablation experiments on the mffw data. This includes the first 10 pairs of partially focused color images under various exposures and obvious misalignment due to severe defocus spread effects. First, the coefficient matrix and the detail layer of quaternion representation of each input color image are extracted using Algorithm 1. Then, base-scale and detail-scale fused results are obtained using the focus measures from Algorithm 2. These results are subsequently combined to generate the final output using the QSSR strategy.
Figure 8: Visual comparison of the background-focused input and various fused results including QBDF-Base-scale, QBDF-Detail-scale and QSSR results. The first row represents the input and fused results. The second row represents the differences between the fused results and corresponding background-focused input.
Fig. 8 provides visual comparison of the background-focused input color image and three fused results. As can be observed, QBDF-Detail-scale fusion introduces noticeable artifacts in the wall regions and creates unrealistic boundary lines within the yellow bounding box. In contrast, the QBDF base-scale fusion achieves smoother integration in low-texture areas but fails to preserve fine details in structured regions such as piano keyboards, resulting in oversmoothed edges. QSSR addresses these problems and significantly improves fusion performance in low-gradient, small-area focused, and boundary regions.
The analysis of the QBDF and the QSSR strategies is also quantitatively presented in this ablation study. Table III gives the quantitative fusion results with different scales of measures. As can be observed, QSSR provides the best quantitative results among five metrics. To verify the robustness of our QSSR strategy, The weight $\tau$ of QSSR is discussed here. There are two different settings for $\tau$ in Eq. (17). The adaptive weight $\tau$ in Eq. (18) provides the best quantitative results among five metrics.
Effectiveness of QMCIF. Table III also evaluates the impact of different representation strategies on fusion performance by comparing our QMCIF framework with conventional real-domain fusion method. To separate the benefits of quaternion representation, we replicate the fusion process in the real domain by converting color images to grayscale. Quantitative results demonstrate that the quaternion representation achieves superior performance across all metrics, particularly outpeforming real domain-based framework by $12\%$ in $Q_{CB}$.
V Conclusion
We proposed a quaternion multi-focus color image fusion framework to perform high-quality color image fusion completely in the quaternion domain. This framework is flexible to be extended from fusing two color images at a time into fusing multiple color images simultaneously under complex challenging scenarios. This framework proposed a quaternion sparse decomposition model for high-precision focus detection of color images. This model jointly learns a low-rank structured coefficient matrix and a detail-scale layer. They both were utilized to estimate the focus level of the image. To solve the optimization problem of the QSD model, we developed an iterative algorithm under the framework of the quaternion alternating direction methods of multipliers. A quaternion base-detail fusion strategy was also introduced for high-quality base-scale and detail-scale color image fusion individually. A quaternion structural similarity strategy was further introduced to balance the trade-off between focus information preservation and artifact removal. Our framework fuses two-scale quaternion matrices and compares them with the input quaternion representations to adaptively select the focused patches with a higher QWSSIM value to form the final fused results. Extensive experiment results have shown that our framework outperforms the state-of-the-art methods in terms of focus information preservation and artifact removal under real-world challenging scenarios.

### v2

@@
-Quaternion Multi-focus Color Image Fusion
+Quaternion Sparse Decomposition for Multi-focus Color Image Fusion
 Abstract
 Multi-focus color image fusion refers to integrating multiple partially focused color images to create a single all-in-focus color image.
-However, existing methods struggle with complex real-world scenarios due to limitations in handling color information and intricate textures.
+However, existing methods struggle with complex real-world scenarios due to limitations in handling color information and intricate focus information.
 To address these challenges, this paper proposes a quaternion multi-focus color image fusion framework to perform high-quality color image fusion completely in the quaternion domain.
-This framework introduces 1) a quaternion sparse decomposition model to jointly learn fine-scale image details and structure information of color images in an iterative fashion for high-precision focus detection, 2) a quaternion base-detail fusion strategy to individually fuse base-scale and detail-scale results across multiple color images for preserving structure and detail information, and 3) a quaternion structural similarity refinement strategy to adaptively select optimal patches from initial fusion results and obtain the final fused result for preserving fine details and ensuring spatially consistent outputs.
+This framework introduces 1) a quaternion consistency-aware focus detection method to jointly learn fine-scale details and structure information of color images and generate patch-wise dual-scale focus maps for high-precision focus detection, 2) a quaternion base-detail fusion strategy to obtain dual-scale initial fusion results across input color images, and 3) a quaternion structural similarity refinement strategy to adaptively select optimal patches from initial fusion results and produce the final fused result that preserves fine details and spatial consistency.
 Extensive experiments demonstrate that the proposed framework outperforms state-of-the-art methods.
 Index Terms:
-Multi-focus color image fusion, quaternion color image fusion, quaternion sparse decomposition.
-publicationid: pubid: $0$–$00$ 2025 IEEE
+Multi-focus color image fusion, quaternion color image fusion, quaternion image decomposition.
+publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
 I Introduction
-A single color image often fails to capture all objects in focus due to optical limitations of cameras [1].
-When imaging scenes with objects at varying distances, only regions within the focal plane are focused while areas outside this plane exhibit defocus blur.
-To circumvent this limitation, multi-focus color image fusion (MCIF) techniques integrate multiple partially focused color images of the same scene using focus detection and adaptive fusion strategies to obtain an all-in-focus color image [2].
-This technology has numerous applications across diverse domains including digital photography, preprocessing operations [3], and target recognition [4, 5].
-Existing MCIF methods can be broadly categorized into deep learning-based and conventional methods.
-Currently, deep learning-based methods have extensively promoted the development of MCIF to enhance the ability of learning multi-scale features to distinguish focused pixels [6, 7].
-Since it is impractical to obtain perfect ground truth (i.e., all-in-focus color images) for real-world scenes [8], many deep learning-based works heavily rely on synthetic training data with a large-scale training process.
-Unlike deep learning-based MCIF methods, conventional methods only depend on human prior knowledge to distinguish focused objects without the training phase [9].
-Existing conventional MCIF methods often treat color image fusion as a direct extension of multi-focus grayscale image fusion techniques by channel-wise processing [10] or grayscale conversion [11].
-For channel-wise processing, these MCIF methods simply treat color channels as three independent grayscale images and directly apply image fusion techniques to each channel to obtain the fused color image.
-For instance, Yang et al.
-[10] extended sparse coding techniques to MCIF by separately measuring focus levels in each channel and fusing sparse coefficients channel-by-channel.
-This may introduce spectral-spatial inconsistencies and color distortions in the fused color images [12, 13].
-For grayscale conversion, existing conventional MCIF methods first convert the input color images into grayscale images and perform intensity-driven focus detection [14, 11] to derive focus maps from grayscale inputs to guide the fusion of original color channels [15].
+A single color image typically fails to maintain all scene objects simultaneously in focus due to inherent optical limitations of camera lenses [1].
+Specifically, when capturing scenes containing objects at varying depths, only regions within the camera’s focal plane appear sharp while areas outside this plane inevitably suffer from defocus blur.
+To overcome this limitation, multi-focus color image fusion (MCIF) techniques integrate multiple partially-focused color images of the same scene by employing focus detection and appropriate fusion strategies to produce a single all-in-focus color image [1, 2].
+This technology has numerous practical applications, such as digital photography enhancement, preprocessing for subsequent image analysis tasks [3], image segmentation [2] and object recognition [4].
+Recent advances in deep learning have significantly promoted MCIF development to enhance multi-scale feature extraction capabilities and improve focused-region discrimination through decision map-based networks [5, 6] and end-to-end architectures [7, 8, 9].
+However, it is impractical to acquire ideal ground truth (all-in-focus color images) in real-world conditions [10].
+These deep-learning-based methods typically rely on synthetic training datasets for training [11].
+Consequently, their generalization capability on real-world images remains limited.
+Conventional MCIF methods depend only on human prior knowledge to identify focused objects without the training phase [12].
+These methods commonly extend grayscale fusion strategies directly to color images either using grayscale conversion [13] or channel-wise processing [14].
+Grayscale-conversion-based MCIF methods first convert the input color images into grayscale images and perform intensity-driven focus detection to obtain focus maps from the grayscale inputs for guiding the fusion process of original color images [13].
 For instance, [15] averaged color channels into a grayscale image and obtained focus maps by multi-scale morphological filtering.
 [16] transformed the color image into the YUV color space and detected focus levels employing the zero crossing and canny edge detectors.
-Although these methods avoid explicit channel separation, they fail to detect multi-channel sharpness variations among the input color images and may produce inaccurate focus maps under complex real-world scenarios.
-This leads to ghosting artifacts or blurred boundaries in the fused results [15, 13].
-Therefore, it remains a challenge to perform high-quality color image fusion in real-world applications.
-The limitations of conventional MCIF methods lie mainly in two key challenges: (1) small-scale focus detection failure: existing techniques struggle to distinguish and retain small focused regions within images containing extensive defocused areas based on grayscale focus detection [4]; (2) boundary artifacts: focus maps derived from intensity-based or edge-based focus detection often produce inaccurate results with visual artifacts introduced particularly in uncertain boundary regions (between focused and defocused areas) and low-gradient regions due to defocus spread effects [15, 18].
-Recently, a quaternion higher-order singular value decomposition (QHOSVD) model [19] was developed to perform quaternion-domain multi-focus color image fusion by fully considering highly-correlated color information.
-QHOSVD treats three color channels as a unified entity of three imaginary parts in a pure quaternion matrix using quaternion representation [20, 21] and extends this representation into a third-order quaternion tensor.
-This effectively preserves cross-channel relationships during color image processing [22, 23, 24].
-Taking advantage of higher-order singular value decomposition in the quaternion domain, QHOSVD extracted focus levels corresponding to each source color image and employed a simple patch-wise fusion strategy to ensure spatial consistency and color fidelity [19].
-However, QHOSVD still fails in focus information preservation and artifact suppression in complex real-world scenarios due to defocus spread effects [18].
-To address these issues, we propose a quaternion multi-focus color image fusion (QMCIF) framework to produce high-quality fused color images. Our main contributions are presented as follows:
-  • We propose a QMCIF framework to perform high-quality color image fusion completely in the quaternion domain.
-QMCIF is flexibly extended from fusing only two color images at a time into fusing multiple color images simultaneously under various challenging scenarios.
-  • To effectively measure the focus level in both small focused areas and uncertain boundary regions, QMCIF introduces a quaternion sparse decomposition model to jointly learn fine-scale image details and structure information of color images in an iterative fashion.
-  • To preserve color structure and detail information, QMCIF introduces a patch-wise quaternion base-detail fusion strategy to fuse multiple partially focused color images in detail-scale and base-scale individually.
-  • To balance the trade-off between focus information preservation and artifact removal, QMCIF further introduces a quaternion structural similarity refinement strategy to flexibly form the final fused results.
-  • Extensive experiments on various datasets demonstrate that QMCIF outperforms the state-of-the-art methods.
+However, these methods may overlook multi-channel sharpness variations and cross-channel correlations.
+This leads to inaccurate focus detection, ghosting artifacts, and blurred boundaries in fused color results [17, 18].
+Channel-wise processing-based MCIF methods treat RGB channels independently and apply grayscale fusion methods separately to each channel.
+For instance, Yang et al.
+[14] applied sparse coding techniques to MCIF and separately measured focus levels in each channel and fused sparse coefficients channel-by-channel.
+They fail to consider the cross-channel correlations and introduce spectral-spatial inconsistencies and undesirable color distortions in the fused results [19].
+Quaternion representation as a promising tool of color image processing can overcome the aforementioned limitations of color image processing in conventional MCIF methods [20, 21].
+It effectively captures color information and preserves cross-channel correlations by treating RGB channels as imaginary components in the quaternion domain [22, 23].
+Recently, a quaternion higher-order singular value decomposition (QHOSVD) model [21] was developed by modeling multi-focus color images as third-order quaternion tensors and successfully maintained cross-channel consistency in the stages of focus detection and fusion.
+However, QHOSVD selects fusion patches based on singular-value energy.
+This may result in spatial inconsistency and degrade the visual quality of the fused color images.
+Motivated by these critical issues, we propose a quaternion multi-focus color image fusion (QMCIF) framework to explicitly designed to simultaneously ensure high-quality multi-focus color image fusion in both spatial and color channels dimensions.
+Our main contributions are presented as follows:
+  • We propose a QMCIF framework to perform color image fusion completely in the quaternion domain. QMCIF is able to obtain high-quality fused color images under various complex scenarios.
+  • To effectively measure the focus level and enhance the capability of focus detection in uncertain regions, QMCIF introduces a quaternion consistency-aware focus detection method to explicitly learn focus features of color images and derive dual-scale focus maps in both high-texture and low-gradient regions.
+  • To preserve color structure and detail information, QMCIF introduces a patch-wise quaternion base-detail fusion strategy to fuse partially focused color images in detail-scale and base-scale individually.
+  • To balance the trade-off between focus information preservation and artifact removal, QMCIF further introduces a quaternion structural similarity refinement strategy to produce the final high-quality fused color image.
+  • Extensive experiments on various challenging scenarios demonstrate that QMCIF outperforms the state-of-the-art methods.
 The rest of this paper is organized as follows: Section II presents the preliminaries of quaternion algebra.
-Section III introduces our framework and extends our framework to multi-color image fusion.
+Section III introduces our framework.
 Section IV presents the experiments and comparisons.
@@
 ${\bm{\mathrm{I}}}_{d}$, $\dot{\bm{\mathrm{I}}}_{d}$ real identity matrix, quaternion identity matrix
-${(\cdot)}^{\mathrm{T}}$$\overline{(\cdot)}$,${(\cdot)}^{\mathrm{H}}$, ${(\cdot)}^{-1}$ transpose, conjugate , conjugate transpose and inverse representation
+[…]
 ${\|\cdot\|}_{1}$, ${\|\cdot\|}_{F}$, $\|\cdot\|$$_{*}$ ${\ell}_{1}$ norm, Frobenius norm, nuclear norm
@@
   4. 4)
-    $Frobenius$ norm: ${\|\dot{\bm{\mathrm{A}}}\|}_{F}={(\sum_{s=1}^{M}\sum_{t=1}^{N}{{|\dot{a}_{s,t}|}^{2}})}^{\frac{1}{2}}$
+    FrobeniusFrobenius norm: ${\|\dot{\bm{\mathrm{A}}}\|}_{F}={(\sum_{s=1}^{M}\sum_{t=1}^{N}{{|\dot{a}_{s,t}|}^{2}})}^{\frac{1}{2}}$
 Definition 2.
@@
 Definition 3.
-Given a color image $\bm{\mathrm{I}}$ in real three-dimensional space, its quaternion representation is defined as:
-$\begin{split}\dot{\bm{\mathrm{I}}}(s,t)=\bm{\mathrm{I}}_{r}(s,t)i+\bm{\mathrm{I}}_{g}(s,t)j+\bm{\mathrm{I}}_{b}(s,t)k\end{split}$
+Given a color image $\bm{\mathrm{I}}$ in three-dimensional real space, its quaternion representation is defined as:
+$\begin{split}\dot{\bm{\mathrm{I}}}(s,t)=\bm{\mathrm{I}}_{r}(s,t)i+\bm{\mathrm{I}}_{g}(s,t)j+\bm{\mathrm{I}}_{b}(s,t)k\end{split}$ (1)
 where $\dot{\bm{\mathrm{I}}}(s,t)$ is the quaternion representation of the color image pixel at the location of $(s,t)$.
@@
 Definition 5.
-Quaternion Structural Similarity (QSSIM) [28] measures the similarity between two color images in the quaternion domain. It is defined by
-$QSSIM({\dot{\bm{\mathrm{X}}}},{\dot{\bm{\mathrm{Y}}}})=\overline{(\dot{a})}\dot{b},$ (1)
+Quaternion Structural Similarity ($Q_{SSIM}$) [27] measures the similarity between two color images in the quaternion domain. It is defined by
+$Q_{SSIM}({\dot{\bm{\mathrm{X}}}},{\dot{\bm{\mathrm{Y}}}})=\overline{(\dot{a})}\dot{b},$ (2)
 where
-$\dot{a}=\frac{2\overline{\dot{\mu}_{\dot{\bm{\mathrm{X}}}}}\dot{\mu}_{\dot{\bm{\mathrm{Y}}}}+C_{1}}{\overline{{\dot{\bm{\mathrm{\mu}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}+C_{1}},$ $\dot{b}=\frac{2\overline{\dot{\sigma}_{\dot{\bm{\mathrm{X}}}}}\dot{\sigma}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}{\overline{{\dot{\bm{\mathrm{\sigma}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}.$ (2)
+$\dot{a}=\frac{2\overline{\dot{\mu}_{\dot{\bm{\mathrm{X}}}}}\dot{\mu}_{\dot{\bm{\mathrm{Y}}}}+C_{1}}{\overline{{\dot{\bm{\mathrm{\mu}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}+C_{1}},$ $\dot{b}=\frac{2\overline{\dot{\sigma}_{\dot{\bm{\mathrm{X}}}}}\dot{\sigma}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}{\overline{{\dot{\bm{\mathrm{\sigma}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}.$
 $\dot{\bm{\mathrm{X}}}$ and $\dot{\bm{\mathrm{Y}}}$ are input patches.
@@
 III Proposed Framework
-This section introduces our quaternion multi-focus color image fusion (QMCIF) framework.
+This section presents our quaternion multi-focus color image fusion (QMCIF) framework in detail.
 Section A provides an overview of our QMCIF framework.
-Section B introduces our quaternion sparse decomposition model and its optimization in detail.
+Section B introduces our quaternion consistency-aware focus detection method.
 Section C presents our quaternion base-detail fusion strategy.
 Section D presents our quaternion structural similarity refinement strategy.
-Section E extends our framework to multiple-color image fusion.
 III-A Overview
-Our QMCIF framework is a patch-wise image fusion framework as shown in Fig.
-1.
-Different from most existing works that perform color image fusion independently, our framework can handle all color channels simultaneously within a single fusion process.
-We first encode the input color images into their quaternion representations and then use the quaternion sparse decomposition model to jointly learn the optimal coefficient matrix and detail layer for each input image for high-precision focus detection.
-We employ the quaternion base-detail fusion strategy to fuse the color images in base and detail scales.
-In the detail-scale fusion process, the two-directional gradients of the detail layers are used to estimate the focus level of each input quaternion representation and select focused patches, generating a detail-scale fused result.
-In the base-scale fusion process, coefficient matrices are combined with the gradients of the corresponding detail layers to assess the focus level of each input quaternion representation and select focused patches to form the base-scale fused result.
-To further improve the fusion quality, we design a quaternion structural similarity refinement strategy to merge these fused results patch-by-patch to produce the final fused result using a quaternion weighted structural similarity index measure.
-Figure 2: Our quaternion sparse decomposition (QSD) model.
-III-B Quaternion sparse decomposition
-III-B1 Model formulation
-Given a partially focused color image $\bm{\mathrm{I}}$ $\in\mathbbm{R}^{M\times N\times 3}$ and its corresponding quaternion representation $\dot{\bm{\mathrm{I}}}$.
-Our goal is to decompose texture information from the static background as a strong focus measure.
-Our quaternion sparse decomposition (QSD) can be formulated as follows.
-$\mathop{\arg\min}_{\dot{\bm{\mathrm{D}}},\dot{\bm{\mathrm{Z}}}}\ \ \sum_{k=1}^{K}{\|\dot{\bm{\mathrm{Z}}}_{k}\|}_{*}+\alpha\sum_{n=1}^{2}\|\nabla_{n}\dot{\bm{\mathrm{B}}}\|_{1}$ (3)
-$+\beta\|\dot{\bm{\mathrm{D}}}\|_{1}+\lambda\|\dot{\bm{\mathrm{E}}}\|_{F}^{2}$
-$s.t.\ \ \dot{\bm{\mathrm{I}}}=\dot{\bm{\mathrm{B}}}+\dot{\bm{\mathrm{D}}}+\dot{\bm{\mathrm{E}}},\ \mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}=\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}_{k}.$
-In Eq.
-(3), ${\nabla}_{1}$ and ${\nabla}_{2}$ represent the vertical and horizontal gradients in the quaternion domain, respectively.
-The operator $\mathcal{R}(\cdot)$ extracts image patches of size $\sqrt{d}\times\sqrt{d}$ and stacks their vectorized forms into a quaternion matrix in $\mathbbm{H}^{d\times P}$.
-Column vectors can be divided into $K$ groups.
-A set of groups is denoted as $\{\mathcal{R}(\dot{\bm{\mathrm{B}}})_{1},\mathcal{R}(\dot{\bm{\mathrm{B}}})_{2},\cdots,\mathcal{R}(\dot{\bm{\mathrm{B}}})_{K}\}$, namely $\mathcal{R}(\dot{\bm{\mathrm{B}}})$ where $\mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}$ represents the $k^{th}$ group.
+Fig.
+1 illustrates our quaternion multi-focus color image fusion (QMCIF) framework.
+Unlike most existing methods that process color images in a grayscale manner, QMCIF performs joint fusion across all color channels within a unified quaternion representation.
+Input color images $\bm{\mathrm{I}}_{1}$, $\bm{\mathrm{I}}_{2}$ are converted into their quaternion representations $\dot{\bm{\mathrm{I}}}_{1}$, $\dot{\bm{\mathrm{I}}}_{2}$ using Eq.
+(1).
+A quaternion consistency-aware focus detection method jointly estimates their optimal coefficient matrices and detail layers to construct dual-scale focus maps for accurate focus estimation.
+Under the guidance of dual-scale focus maps, a patch-wise quaternion fusion strategy is applied to fuse the quaternion representations $\dot{\bm{\mathrm{I}}}_{1}$ and $\dot{\bm{\mathrm{I}}}_{2}$ at the base and detail scales to generate the initial fusion results $\dot{\bm{\mathrm{F}}}_{1}$ and $\dot{\bm{\mathrm{F}}}_{2}$.
+They are further applied with a quaternion structure-similarity-based refinement strategy patch-by-patch to generate the final fused result $\dot{\bm{\mathrm{F}}}$ that is converted back into the real domain to obtain a high-quality all-in-focus color image.
+Notably, in additional to fuse two color images at a time, our QMCIF framework can be applied to fuse multiple input color images simultaneously.
+This advantage is verified by the experiments in Figs.
+7 and 8.
+III-B Quaternion consistency-aware focus detection
+To achieve high-precision focus estimation in both high-texture and low-gradient regions, our quaternion consistency-aware focus detection method contains two main steps: (1) a quaternion focal element decomposition model to effectively extract focus-related features in the quaternion domain; and (2) a patch-wise dual-scale focus map generation strategy.
+Its flowchart is shown in Fig.
+2.
+Figure 2: Our quaternion consistency-aware focus detection (QCAFD).
+III-B1 Quaternion focal element decomposition
+Given an input quaternion representation $\dot{\bm{\mathrm{I}}}$ $\in\mathbbm{H}^{M\times N}$, our quaternion focal element decomposition (QFED) model decomposes texture information from its static background as a strong focus measure.
+Our QFED can be formulated as follows.
+$\mathop{\arg\min}_{\dot{\bm{\mathrm{D}}},\dot{\bm{\mathrm{Z}}}}$ $\sum_{k=1}^{K}{\|\dot{\bm{\mathrm{Z}}}_{k}\|}_{*}+\alpha({\|\nabla_{1}\dot{\bm{\mathrm{B}}}\|}_{1}+{\|\nabla_{2}\dot{\bm{\mathrm{B}}}\|}_{1})$ (3)
+$+\beta\|\dot{\bm{\mathrm{D}}}\|_{1}+\lambda\|\dot{\bm{\mathrm{E}}}\|_{F}^{2},$
+$s.t.\ \ \dot{\bm{\mathrm{I}}}=\dot{\bm{\mathrm{B}}}+\dot{\bm{\mathrm{D}}}+\dot{\bm{\mathrm{E}}},\ \mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}=\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}_{k}$
+where ${\nabla}_{1}$ and ${\nabla}_{2}$ represent the quaternion vertical and horizontal gradients respectively.
+Operator $\mathcal{R}(\cdot)$ extracts image patches of size $\sqrt{d}\times\sqrt{d}$ and stacks their vectorized forms into a quaternion matrix in $\mathbbm{H}^{d\times P}$.
+Column vectors are divided into $K$ groups denoted as $\mathcal{R}(\dot{\bm{\mathrm{B}}})=\{\mathcal{R}(\dot{\bm{\mathrm{B}}})_{1},\mathcal{R}(\dot{\bm{\mathrm{B}}})_{2},\cdots,\mathcal{R}(\dot{\bm{\mathrm{B}}})_{K}\}$ where $\mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}$ represents the $k^{th}$ group.
 $\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}}$ captures the low-rank property of $\mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}$.
-The nuclear norm of $\dot{\bm{\mathrm{Z}}}{{}_{k}}$ serves as a convex relaxation of the rank function.
+$\|\dot{\bm{\mathrm{Z}}}_{k}\|$$_{*}$ serves as a convex relaxation of the rank function.
 The positive parameters $\alpha$, $\beta$ and $\lambda$ balance the effects of each term in the optimization process.
-As displayed in Fig.
-2, our QSD converts a color image to its quaternion representation $\dot{\bm{\mathrm{I}}}$ that is then decomposed into a base-scale layer $\dot{\bm{\mathrm{B}}}$ with low-rank structural information, a detail-scale layer $\dot{\bm{\mathrm{D}}}$ with sparse details and a Gaussian noise layer $\dot{\bm{\mathrm{E}}}$.
-All layers are quaternion matrices belonging to $\mathbbm{H}^{M\times N}$.
-To fully utilize the focus information in $\dot{\bm{\mathrm{B}}}$, QSD uses nonlocal similarity to extract local patches in $\dot{\bm{\mathrm{B}}}$, stack these patches into column vectors, and partition them into several groups using the K-means clustering method.
-QSD learns a set of coefficient matrices $\dot{\bm{\mathrm{Z}}}_{k}$ from these groups using the quaternion dictionary $\dot{\bm{\mathrm{A}}}$ $\in\mathbbm{H}^{d\times L}$ containing $L$ atoms and concatenates these coefficient matrices to obtain $\dot{\bm{\mathrm{Z}}}$ $\in\mathbbm{H}^{L\times P}$.
+Our QFED decomposes $\dot{\bm{\mathrm{I}}}$ into a base-scale layer $\dot{\bm{\mathrm{B}}}$ with low-rank structural information, a detail-scale layer $\dot{\bm{\mathrm{D}}}$ with sparse details and a Gaussian noise layer $\dot{\bm{\mathrm{E}}}$.
+To fully utilize the focus information in $\dot{\bm{\mathrm{B}}}$, QFED uses nonlocal similarity to extract local patches in $\dot{\bm{\mathrm{B}}}$, stacks these patches into column vectors, and partitions them into several groups using K-means clustering method.
+QFED learns a set of coefficient matrices $\dot{\bm{\mathrm{Z}}}_{k}$ from these groups using pretrained quaternion dictionary $\dot{\bm{\mathrm{A}}}$ $\in\mathbbm{H}^{d\times L}$ and concatenates these coefficient matrices to obtain $\dot{\bm{\mathrm{Z}}}$ $\in\mathbbm{H}^{L\times P}$.
 This ensures that patches within the same group are similar.
-In general, the detail layer intuitively reflects the sharpness of edge pixels, as it exhibits a greater amplitude of pixel intensity in focused regions.
-However, in low-texture and boundary areas, the detail layer $\dot{\bm{\mathrm{D}}}$ often provides insufficient features.
-This may degrade fusion performance.
-To capture the focus information of smooth regions, coefficient matrix $\dot{\bm{\mathrm{Z}}}$ is introduced as a base-scale measure using quaternion sparse coding [21].
-Obtaining the optimal $\dot{\bm{\mathrm{D}}}$ and $\dot{\bm{\mathrm{Z}}}$ simultaneously offers our QSD the greater flexibility than handcrafted filters.
-III-B2 Optimization
-To solve the optimization problem of our QSD in Eq.
-(3), we here introduce an iterative learning algorithm under the framework of the quaternion alternating direction method of multipliers [25, 29].
-According to the definitions and lemmas in Section II, we write the Lagrangian function of our QSD in Eq.
+Optimization.
+To solve the optimization problem of our QFED in Eq.
+(3), we here introduce an iterative learning algorithm under the framework of the quaternion alternating direction method of multipliers [24].
+According to the definitions and lemmas in Section II, we write the Lagrangian function of our QFED in Eq.
 (3) as follows:
@@
 $+\frac{\mu}{2}\|\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}-\dot{\bm{\mathrm{D}}}-\dot{\bm{\mathrm{E}}}\|_{F}^{2}.$
-[…]
+where ${\{{\dot{\bm{\mathrm{J}}}}_{k}\}}_{k=1}^{K}\in\mathbbm{H}^{L\times P}$ represent a set of auxiliary variables to replace ${\{{\dot{\bm{\mathrm{Z}}}}_{k}\}}_{k=1}^{K}$.
+Matrices ${\dot{\bm{\mathrm{G}}}}_{1}$ and ${\dot{\bm{\mathrm{G}}}}_{2}\in\mathbbm{H}^{M\times N}$ correspond to quaternion horizontal and vertical gradient matrices, replacing $\nabla_{1}\dot{\bm{\mathrm{B}}}$ and $\nabla_{2}\dot{\bm{\mathrm{B}}}$ respectively.
+Variables ${\{{\dot{\bm{\mathrm{Y}}}}_{1,k}\}}_{k=1}^{K}$ $\in\mathbbm{H}^{L\times P}$, ${\{{\dot{\bm{\mathrm{Y}}}}_{2,k}\}}_{k=1}^{K}$ $\in\mathbbm{H}^{d\times P}$, ${\dot{\bm{\mathrm{Y}}}}_{3},{\dot{\bm{\mathrm{Y}}}}_{4}$ and ${\dot{\bm{\mathrm{Y}}}}_{5}\in\mathbbm{H}^{M\times N}$ are the quaternion Lagrangian multipliers in the optimization process.
+$\left\langle\cdot\right\rangle$ denotes a quaternion trace product while $\mu$ is a penalty factor.
+Variables $\dot{\bm{\mathrm{J}}}{{}_{k}}$, $\dot{\bm{\mathrm{Z}}}{{}_{k}}$, $\dot{\bm{\mathrm{Y}}}{{}_{1,k}}$ and $\dot{\bm{\mathrm{Y}}}{{}_{2,k}}$, $\dot{\bm{\mathrm{G}}}{{}_{1}}$, $\dot{\bm{\mathrm{G}}}{{}_{2}}$, $\dot{\bm{\mathrm{D}}}$, $\dot{\bm{\mathrm{E}}}$, $\dot{\bm{\mathrm{Y}}}{{}_{3}}$, $\dot{\bm{\mathrm{Y}}}{{}_{4}}$, $\dot{\bm{\mathrm{Y}}}{{}_{5}}$ are updated in each iteration.
 Update ${\{\dot{\bm{\mathrm{J}}}{{}_{k}}\}}_{k=1}^{K}$.
@@
 Let $\dot{\bm{\mathrm{P}}}=$ $\dot{\bm{\mathrm{Z}}}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}}{\mu}$.
-$\dot{\bm{\mathrm{J}}}{{}_{k}}$ is minimized using the quaternion singular value thresholding method [26].
-$\dot{\bm{\mathrm{J}}}{{}_{k}}$ $=\mathop{\arg\min}_{\dot{\bm{\mathrm{J}}}{{}_{k}}}\ \ {\|\dot{\bm{\mathrm{J}}}_{k}\|}_{*}+\frac{\mu}{2}\|\dot{\bm{\mathrm{J}}}{{}_{k}}-(\dot{\bm{\mathrm{Z}}}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}}{\mu})\|_{F}^{2}$ (5)
-$=\mathop{\arg\min}_{\dot{\bm{\mathrm{J}}}{{}_{k}}}\ \ \frac{1}{\mu}{\|\dot{\bm{\mathrm{J}}}_{k}\|}_{*}+\frac{1}{2}\|\dot{\bm{\mathrm{J}}}{{}_{k}}-\dot{\bm{\mathrm{P}}}\|_{F}^{2}.$
-Therefore, Eq. (5) can be solved using the Lemma 1.
+The subproblem of $\dot{\bm{\mathrm{J}}}{{}_{k}}$ is reduced to:
+$\dot{\bm{\mathrm{J}}}{{}_{k}}=\mathop{\arg\min}_{\dot{\bm{\mathrm{J}}}{{}_{k}}}\ \ \frac{1}{\mu}{\|\dot{\bm{\mathrm{J}}}_{k}\|}_{*}+\frac{1}{2}\|\dot{\bm{\mathrm{J}}}{{}_{k}}-\dot{\bm{\mathrm{P}}}\|_{F}^{2}.$ (5)
+Eq. (5) is solved using the Lemma 1 in Section II.
 Update ${\{\dot{\bm{\mathrm{Z}}}{{}_{k}}\}}_{k=1}^{K}$.
@@
 $\mathop{\arg\min}_{\dot{\bm{\mathrm{Z}}}{{}_{k}}}\ \ \frac{\mu}{2}{\|\dot{\bm{\mathrm{Z}}}{{}_{k}}-\dot{\bm{\mathrm{Q}}}\|}_{F}^{2}+\frac{\mu}{2}\|\mathcal{R}(\dot{\bm{\mathrm{B}}}){{}_{k}}-\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{2,k}}}{\mu}\|_{F}^{2}.$ (6)
-We set the gradient of Eq. (6) with respect to $\dot{\bm{\mathrm{Z}}}{{}_{k}}$ to zero. $\dot{\bm{\mathrm{Z}}}{{}_{k}}$ is explicitly updated as follows.
+$\dot{\bm{\mathrm{Z}}}{{}_{k}}$ is explicitly updated as follows.
 $\begin{split}\dot{\bm{\mathrm{Z}}}{{}_{k}}={(\dot{\bm{\mathrm{A}}}{{}^{\mathrm{H}}}\dot{\bm{\mathrm{A}}}+\dot{\bm{\mathrm{I}}}_{d})}^{-1}(\dot{\bm{\mathrm{A}}}{{}^{\mathrm{H}}}({\mathcal{R}(\dot{\bm{\mathrm{B}}})}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{2,k}}}{\mu})+\dot{\bm{\mathrm{Q}}}).\end{split}$ (7)
@@
 $\dot{\bm{\mathrm{G}}}{{}_{1}}$ is updated according to the subproblem:
-[…]
+$\begin{aligned} \dot{\bm{\mathrm{G}}}{{}_{1}}=\mathop{\arg\min}_{\dot{\bm{\mathrm{G}}}{{}_{1}}}\ \ \frac{\alpha}{\mu}{\|{\dot{\bm{\mathrm{G}}}}_{1}\|}_{1}+\frac{1}{2}\|\dot{\bm{\mathrm{G}}}{{}_{1}}-\dot{\bm{\mathrm{M}}}_{1}\|{{}_{F}^{2}}\end{aligned}$ (8)
 The solution to Eq.
-(8) is obtained using the soft-thresholding method in Lemma 2.
+(8) is obtained using the soft-thresholding method in Lemma 2 in Section II.
 Similarly, $\dot{\bm{\mathrm{G}}}{{}_{2}}$ is updated in the same way as the $\dot{\bm{\mathrm{G}}}{{}_{1}}$-subproblem.
@@
 $\dot{\bm{\mathrm{B}}}$ is updated as below.
-$\begin{split}\dot{\bm{\mathrm{B}}}=\mathcal{F}{{}^{-1}}({\frac{\dot{\bm{\mathrm{\Sigma}}}}{\mathcal{F}(\nabla_{1}^{\mathrm{T}}\nabla_{1}+\nabla_{2}^{\mathrm{T}}\nabla_{2})+2}}).\end{split}$ (10)
+[…]
 Update $\dot{\bm{\mathrm{D}}}$.
@@
 $\begin{split}\dot{\bm{\mathrm{D}}}=\mathop{\arg\min}_{\dot{\bm{\mathrm{D}}}}\ \ \frac{\beta}{\mu}{\|\dot{\bm{\mathrm{D}}}\|}_{1}+\frac{1}{2}\|\dot{\bm{\mathrm{D}}}-\dot{\bm{\mathrm{M}}}_{7}\|_{F}^{2}.\end{split}$ (11)
-Eq. (11) can also be solved using Lemma 2.
+Eq. (11) can also be solved using Lemma 2 in Section II.
 Update $\dot{\bm{\mathrm{E}}}$.
@@
 $\dot{\bm{\mathrm{E}}}={(2\lambda+\mu\dot{\bm{\mathrm{I}}}_{d})}^{-1}(\mu\dot{\bm{\mathrm{M}}}_{8}),$ (12)
-The optimization process of the QSD model is shown in Algorithm 1.
+The optimization process of the QFED model is shown in Algorithm 1.
 Input: The quaternion representation of the source image $\dot{\bm{\mathrm{I}}}$, the parameters $\mu$, $\lambda$, $\alpha$ and $\beta$, the iteration number $t$.
 Output: Optimal quaternion coefficient matrix $\dot{\bm{\mathrm{Z}}}$ and detail layer $\dot{\bm{\mathrm{D}}}$
-1 Initialize ${\{\dot{\bm{\mathrm{J}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$,${\{\dot{\bm{\mathrm{Z}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$, ${\{\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{0}}\}}{{}_{k=1}^{K}}$,${\{\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{0}}\}}{{}_{k=1}^{K}}$, $\dot{\bm{\mathrm{G}}}{{}_{1}^{0}}$, $\dot{\bm{\mathrm{G}}}{{}_{2}^{0}}$, $\dot{\bm{\mathrm{B}}}{{}^{0}}$, $\dot{\bm{\mathrm{D}}}{{}^{0}}$, $\dot{\bm{\mathrm{E}}}{{}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{3}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{4}^{0}}$ and $\dot{\bm{\mathrm{Y}}}{{}_{5}^{0}}$;
+1 Initialize ${\{\dot{\bm{\mathrm{J}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$${\{\dot{\bm{\mathrm{Z}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$, ${\{\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{0}}\}}{{}_{k=1}^{K}}$${\{\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{0}}\}}{{}_{k=1}^{K}}$, $\dot{\bm{\mathrm{G}}}{{}_{1}^{0}}$, $\dot{\bm{\mathrm{G}}}{{}_{2}^{0}}$, $\dot{\bm{\mathrm{B}}}{{}^{0}}$, $\dot{\bm{\mathrm{D}}}{{}^{0}}$, $\dot{\bm{\mathrm{E}}}{{}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{3}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{4}^{0}}$ and $\dot{\bm{\mathrm{Y}}}{{}_{5}^{0}}$;
 2while not converged do
@@
 11 for $k=1,\cdots,K$ do
-12             $\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{t}}=\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{t-1}}+\mu(\dot{\bm{\mathrm{Z}}}{{}_{k}^{t}}-\dot{\bm{\mathrm{J}}}{{}_{k}^{t}})$;
-13            $\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{t}}=\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{t-1}}+\mu(\mathcal{R}(\dot{\bm{\mathrm{B}}}){{}_{k}^{t}}-\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}^{t}})$;
+[…]
+[…]
 14   end for
 15
-16      ${\dot{\bm{\mathrm{Y}}}{{}_{3}^{t}}}={\dot{\bm{\mathrm{Y}}}{{}_{3}^{t-1}}}+\mu(\dot{\bm{\mathrm{G}}}{{}_{1}^{t}}-\nabla_{1}\dot{\bm{\mathrm{B}}}{{}^{t}})$;
-17      ${\dot{\bm{\mathrm{Y}}}{{}_{4}^{t}}}={\dot{\bm{\mathrm{Y}}}{{}_{4}^{t-1}}}+\mu(\dot{\bm{\mathrm{G}}}{{}_{2}^{t}}-\nabla_{2}\dot{\bm{\mathrm{B}}}{{}^{t}})$;
-18      ${\dot{\bm{\mathrm{Y}}}{{}_{5}^{t}}}={\dot{\bm{\mathrm{Y}}}{{}_{5}^{t-1}}}+\mu(\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}{{}^{t}}-\dot{\bm{\mathrm{D}}}{{}^{t}}-\dot{\bm{\mathrm{E}}}{{}^{t}})$;
+[…]
+[…]
+[…]
 19 $\mu=\min{\{{10}^{6},\mu*1.1\}}$;
-20      $t$longleftarrow $t+1$;
+20 $t$ $\longleftarrow$ $t+1$;
 21 end while
-Algorithm 1 Quaternion sparse decomposition
+Algorithm 1 Quaternion focal element decomposition (QFED)
+III-B2 Patch-wise dual-scale focus map generation
+As illustrated in Fig.
+2, we obtain the corresponding detail layers $\dot{\bm{\mathrm{D}}}{{}_{1}}$, $\dot{\bm{\mathrm{D}}}{{}_{2}}$ and coefficient matrices $\dot{\bm{\mathrm{Z}}}{{}_{1}}$, $\dot{\bm{\mathrm{Z}}}{{}_{2}}$ from the QFED process for given input quaternion representations $\dot{\bm{\mathrm{I}}}{{}_{1}}$, $\dot{\bm{\mathrm{I}}}{{}_{2}}$.
+To fully leverage the complementary information in $\dot{\bm{\mathrm{D}}}$ and $\dot{\bm{\mathrm{Z}}}$, we propose a patch-wise dual-scale focus map generation strategy to generate the base-scale and detail-scale focus maps at the patch level.
+Detail Amplification. The detail-scale layer for each pixel location is amplified within a sliding window to strengthen the fine-grained details:
+$\dot{\bm{\mathrm{D}}}{{}_{s}}(x,y)=\sum_{s=-r}^{r}\sum_{t=-r}^{r}\dot{\bm{\mathrm{D}}}(x+s,y+t),$ (13)
+[…]
+Dual-scale focus measure. The dual-scale focus measures used to compute patch-wise focus levels are defined as:
+$l_{B}={\|\nabla_{1}\dot{\bm{\mathrm{d}}}\|}_{1}+{\|\nabla_{2}\dot{\bm{\mathrm{d}}}\|}_{1}+\theta{\|\dot{\bm{\mathrm{z}}}\|}_{2},$ (14)
+$l_{D}=\phi({\|\nabla_{1}\dot{\bm{\mathrm{d}}}_{s}\|}_{1}+{\|\nabla_{2}\dot{\bm{\mathrm{d}}}_{s}\|}_{1}),$ (15)
+where $l_{B}$ and $l_{D}$ represent patch-wise base-scale and detail-scale focus levels of input quaternion representation $\dot{\bm{\mathrm{I}}}$; $\dot{\bm{\mathrm{d}}}$ and $\dot{\bm{\mathrm{z}}}$ denote the local patch and corresponding column vector of $\dot{\bm{\mathrm{D}}}$ and $\dot{\bm{\mathrm{Z}}}$ respectively; $\dot{\bm{\mathrm{d}}}_{s}$ represents the local patch of $\dot{\bm{\mathrm{D}}}_{s}$ computed by Eq.
+(13).
+Parameter $\theta$ is set to 1 by default.
+$\phi(\cdot)$ is a Laplacian-based enhancement function defined as $\phi(x)=1-e^{-x/\gamma}$ where $\gamma=0.2$ by default.
+The detail-scale focus measure captures the magnitude of two-directional gradients that effectively reflect focus levels in color images.
+The base-scale focus measure jointly considers gradient magnitudes and coefficient matrix energy, thereby improving spatial coherence and mitigating local artifacts.
+To compare base-scale and detail-scale focus levels of the local patches in $\dot{\bm{\mathrm{I}}}{{}_{1}}$ and $\dot{\bm{\mathrm{I}}}{{}_{2}}$, the patch-wise dual-scale focus maps $\bm{\mathrm{m}}{{}_{B}}$ and $\bm{\mathrm{m}}{{}_{D}}$ are computed as:
+$\bm{\mathrm{m}}{{}_{B}}=\begin{cases}0,&{\rm if}\ l_{B,1}>l_{B,2}\\ 1,&{\rm if}\ l_{B,1}\leq l_{B,2}\end{cases}$ (16)
+$\bm{\mathrm{m}}{{}_{D}}=\begin{cases}0,&{\rm if}\ l_{D,1}>l_{D,2}\\ 1,&{\rm if}\ l_{D,1}\leq l_{D,2}\end{cases}$ (17)
+where $l_{B,1}$ and $l_{B,2}$ are computed using Eq.
+(14); $l_{D,1}$ and $l_{D,2}$ are computed using Eq.
+(15); $\bm{\mathrm{m}}{{}_{B}}$ and $\bm{\mathrm{m}}{{}_{D}}$ denote patch-wise base-scale and detail-scale focus maps respectively.
+Aggregating these decisions across all patches yields the global base-scale and detail-scale focus maps $\bm{\mathrm{M}}{{}_{B}}$ and $\bm{\mathrm{M}}{{}_{D}}$.
+Figure 3: Our quaternion base-detail fusion strategy.
 III-C Quaternion base-detail fusion strategy
-This subsection proposes a quaternion base-detail fusion (QBDF) strategy to achieve high-quality color image fusion in base and detail scales.
-Fig.
-3 illustrates our QBDF strategy.
-Given quaternion representations ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{2}$ $\in\mathbbm{H}^{2\times M\times N}$ of two input color images, our QBDF strategy first uses the QSD model to learn their detail-scale layers ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{2}$ and coefficient matrices ${\{\dot{\bm{\mathrm{Z}}}^{(i)}\}}_{i=1}^{2}$, and then performs two-scale image fusion.
-To fully utilize the complementary features of coefficient matrices and detail-scale layers, QBDF strategy introduces a two-scale fusion process to obtain base-scale and detail-scale fusion results at the patch level.
-The patch-wise fusion rule is adopted to reduce the sensitivity to the undetermined pixels directly generated by focus detection.
-First, we define two-scale focus measures as below.
-$l_{b,j}=\mathop{\arg\max}_{i}\ \ \sum_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}_{1}+\theta{\|\dot{\bm{\mathrm{Z}}}_{j}^{(i)}\|}_{2},$ (13)
-$l_{d,j}=\mathop{\arg\max}_{i}\ \ \sum_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}_{1},$ (14)
-where $l_{b}$ and $l_{d}$ denote the patch-wise focus maps of base-scale and detail-scale fused results respectively, $j$ defines the center spatial location of each patch and parameter $\theta$ is here set to 1 by default.
-$\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}$ represents the $j^{th}$ patch of $\dot{\bm{\mathrm{D}}}^{(i)}$ and $\dot{\bm{\mathrm{Z}}}{{}^{(i)}_{j}}$ corresponds to the $j^{th}$ column vector of $\dot{\bm{\mathrm{Z}}}^{(i)}$.
-The detail-scale focus measure uses the two-directional gradients of the detail-scale layers.
-The gradients that intuitively represent variations in the quaternion pixels can substantially reflect the focus levels of color images.
-Based on the greater amplitude of Eq.
-(14), $l_{d,j}$ is assigned with the corresponding $i$ in $j^{th}$ patch.
-Figure 3: Flowchart of our quaternion base-detail fusion (QBDF) and quaternion structural similarity refinement (QSSR) strategies.
-Although the detail-scale focus map $l_{d}$ effectively corrects the inaccurate focus map using the neighboring information for boundary and small focused regions, it may produce visual artifacts and fail to handle low-gradient background areas.
-This limitation can lead to significant inconsistencies in the detail-scale fused result.
-To address this issue, the base-scale focus measure integrates the gradients of the detail layers with their corresponding coefficient matrices.
-This ensures local consistency and mitigates spatial artifacts caused by the detail-scale focus map.
-Based on the greater amplitude of Eq.
-(13), $l_{b,j}$ is assigned with the corresponding $i$ in $j^{th}$ patch.
-The focused patches are obtained and aggregated into detail-scale and the base-scale fused results ${\dot{\bm{\mathrm{P}}}}^{(1)}$ and ${\dot{\bm{\mathrm{P}}}}^{(2)}$ using the following two-scale patch-wise fusion rule:
-$\dot{\bm{\mathrm{P}}}{{}^{(1)}_{j}}=\left\{\begin{aligned} &\dot{\bm{\mathrm{I}}}{{}^{(1)}_{j}},&&l_{d,j}=1\\ &\dot{\bm{\mathrm{I}}}{{}^{(2)}_{j}},&&l_{d,j}=2\end{aligned}\right.$ (15)
-$\dot{\bm{\mathrm{P}}}{{}^{(2)}_{j}}=\left\{\begin{aligned} &\dot{\bm{\mathrm{I}}}{{}^{(1)}_{j}},&&l_{b,j}=1\\ &\dot{\bm{\mathrm{I}}}{{}^{(2)}_{j}},&&l_{b,j}=2\end{aligned}\right.$ (16)
+Our quaternion base-detail fusion (QBDF) strategy aims to achieve high-quality multi-focus color image fusion that operates at base and detail scales.
+The overall fusion process is illustrated in Fig.
+3.
+Given the input quaternion representations $\dot{\bm{\mathrm{I}}}_{1}$ $\in\mathbbm{H}^{M\times N}$ and $\dot{\bm{\mathrm{I}}}_{2}$ $\in\mathbbm{H}^{M\times N}$ and corresponding dual-scale focus maps $\bm{\mathrm{M}}_{B}$ and $\bm{\mathrm{M}}_{D}$, the QBDF strategy performs dual-scale image fusion.
+The patch-wise selection results that are derived from the focus maps are used to construct the fused images at each scale.
+Specifically, the base-scale fused result ${\dot{\bm{\mathrm{F}}}}_{1}$ and detail-scale fused result ${\dot{\bm{\mathrm{F}}}}_{2}$ are computed using the following rules:
+${\dot{\bm{\mathrm{F}}}}_{1}=\left\{\begin{aligned} &{\dot{\bm{\mathrm{I}}}}_{1},&&{\rm if}\ \bm{\mathrm{M}}_{B}=0\\ &{\dot{\bm{\mathrm{I}}}}_{2},&&{\rm if}\ \bm{\mathrm{M}}_{B}=1\end{aligned}\right$.
+(18)
+${\dot{\bm{\mathrm{F}}}}_{2}=\left\{\begin{aligned} &{\dot{\bm{\mathrm{I}}}}_{1},&&{\rm if}\ \bm{\mathrm{M}}_{D}=0\\ &{\dot{\bm{\mathrm{I}}}}_{2},&&{\rm if}\ \bm{\mathrm{M}}_{D}=1\end{aligned}\right$.
+(19)
+This fusion rule ensures that the source image with the most focused content is selected and integrated at detail or base scales for each patch location.
 III-D Quaternion structural similarity refinement strategy
-To obtain a high-quality fused result, we introduce a quaternion structural similarity refinement (QSSR) strategy to refine the base-scale and detail-scale fused patches of ${\{\dot{\bm{\mathrm{P}}}^{(i)}\}}_{i=1}^{2}$ using a selection rule with a maximum quaternion image structural similarity value.
-We assume that the patch to be selected should be most similar to the corresponding region of the source image with the largest focus information.
-Therefore, we propose a quaternion weighted structural similarity (QWSSIM) strategy to refine the distorted patches that are wrongly detected as focused ones based on the theory of Definition 5.
-QWSSIM is defined in the following equation.
-$QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)})=\sum_{i=1}^{2}{\tau}_{i,j}QSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)},{\dot{\bm{\mathrm{I}}}}_{j}^{(i)}),$ (17)
-$QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)})=\sum_{i=1}^{2}{\tau}_{i,j}QSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)},{\dot{\bm{\mathrm{I}}}}_{j}^{(i)}),$
-where $j\in[1,P]$ denotes the center spatial location of each patch where $P$ denotes the number of overlapping patches.
-${\tau}_{i,j}$ is an assigned weight to $j^{th}$ quaternion patch of $i^{th}$ input quaternion representation.
-The patches selected from the two-scale fused results by our QBDF strategy should be quite similar to the corresponding focused patches of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}^{2}_{i=1}$.
-In other words, each local quaternion patch of the final fused result has a maximum QWSSIM value with the corresponding region of ${\{\dot{\bm{\mathrm{I}}}\}}^{2}_{i=1}$.
-${\tau}_{i,j}$ is significantly larger when the corresponding region of ${\dot{\bm{\mathrm{I}}}}^{(i)}$ is focused.
-Fig.
-3 illustrates the flowchart of our QSSR strategy.
-Our QSSR strategy further compares two-scale fused results with the input quaternion representations patch-by-patch to assess the patch-wise structural similarity using our QWSSIM measure.
-Patches with higher QWSSIM values are selected to form the final fused result.
-To adaptively compute ${\tau}_{i,j}$, we utilize the detail-scale measure to generate each weight map with the size of $n\times P$ where n depends on the number of inputs.
-The detail-scale measure is enhanced using a Laplacian function $\phi(x)=1-e^{-\frac{x}{\gamma}}$ where $\gamma$ is $=0.2$ by default.
-Then the weight map is formulated in the following equation.
-${\tau}_{i,j}=\frac{\phi({({\sum}_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}{{}_{1}})})}{\sum_{i}\phi(({\sum}_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}{{}_{1}}))}.$ (18)
-We apply a maximum selection rule based on the QWSSIM measure at the patch level.
-It effectively corrects defocused regions in the two-scale fused images.
-Additionally, this strategy eliminates visual artifacts caused by inaccurate focus maps.
-The selection rule is defined as follows.
-${\bm{\mathrm{W}}}_{j}=\left\{\begin{aligned} &1,&QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)})>QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)})\\ &0,&QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)})\leq QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)}).\end{aligned}\right.$ (19)
-The weight matrix $\bm{\mathrm{W}}$ denotes a refined patch-wise focus map.
-When ${\bm{\mathrm{W}}}_{j}=1$, we select the $j^{th}$ patch from ${\dot{\bm{\mathrm{P}}}}^{(1)}$.
-Otherwise, we choose the $j^{th}$ patch from ${\dot{\bm{\mathrm{P}}}}^{(2)}$.
-The final fused result is obtained at the patch level as below:
-$\dot{\bm{\mathrm{F}}}_{j}=\left\{\begin{aligned} &{\dot{\bm{\mathrm{P}}}}_{j}^{(1)},&{\bm{\mathrm{W}}}_{j}=1\\ &{\dot{\bm{\mathrm{P}}}}_{j}^{(2)},&{\bm{\mathrm{W}}}_{j}=0.\end{aligned}\right.$ (20)
-The patches are aggregated to form a quaternion fused result ${\dot{\bm{\mathrm{F}}}}$.
-After the inverse transformation from the quaternion domain into the real domain, we obtain a fused color image $\bm{\mathrm{F}}$ $\in\mathbbm{R}^{M\times N\times 3}$.
-III-E Extension to multiple color image fusion
-Fig.
-1 displays our quaternion multi-focus color image fusion framework to fuse two partially focused color images.
-Our framework is flexible to be extended into fusing multiple color images.
-Let ${\{{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ denote $n$ input color images ($n\geq 2$).
-The multi-focus color image fusion performed on ${\{{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ is decribed in the following process.
-Focus detection.
-We first encode multiple color images into their quaternion representations ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$.
-The QSD model is then used to decompose them into quaternion coefficient matrices ${\{\dot{\bm{\mathrm{Z}}}^{(i)}\}}_{i=1}^{n}$ and quaternion detail layers ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ in the iterative learning process.
-QBDF strategy.
-Our QBDF strategy uses two-scale focus measures to obtain the fused results.
-The detail-scale focus measure uses two-directional gradients of $\dot{\bm{\mathrm{D}}}^{(i)}$ while the base-scale focus measure integrates the gradients of $\dot{\bm{\mathrm{D}}}^{(i)}$ with their corresponding $\dot{\bm{\mathrm{Z}}}^{(i)}$ for each quaternion representation.
-The base-scale and detail-scale measures are easily extended to ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ ($n\geq 2$) using Eqs.
-(13) and (14).
-We select focused patches to obtain base-scale and detail-scale fused results ${\{\dot{\bm{\mathrm{P}}}^{(i)}\}}_{i=1}^{2}$ based on Eqs.
-(15) and (16) similarly.
-QSSR strategy.
-Our QSSR strategy compares each patch from base-scale and detail-scale fused results with input quaternion representations respectively using the QWSSIM.
-The patches with a higher QWSSIM value are selected as focused ones to form a final fused result.
-Therefore, the QBDF and QSSR strategies are easily extended to multiple color images.
-The detailed process is presented in Algorithm 2.
-Input: Input $n$ color images ${\{\bm{\mathrm{I}}^{(i)}\}}{{}_{i=1}^{n}}$ ($n\geq 2$)
-Output: The fused image $\bm{\mathrm{F}}$
-1 for $i=1,\cdots,n$ do
-2       Transform $\bm{\mathrm{I}}^{(i)}\in\mathbbm{R}$ into $\dot{\bm{\mathrm{I}}}^{(i)}$ in the quaternion domain;
-3      Solve the quaternion sparse coefficient matrix ${\dot{\bm{\mathrm{Z}}}}^{(i)}$ and the quaternion detail layer ${\dot{\bm{\mathrm{D}}}}^{(i)}$ using Algorithm 1;
-4 end for
-5Divide each quaternion representation of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ and ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ into $P_{1}$ overlapping patches respectively;
-6Fuse the detail-scale result ${\dot{\bm{\mathrm{P}}}}^{(1)}$ using Eq. (14) ;
-7Divide each quaternion representation of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ and ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ into $P_{2}$ overlapping patches respectively;
-8Fuse the base-scale result ${\dot{\bm{\mathrm{P}}}}^{(2)}$ using Eq. (13);
-9Divide ${\dot{\bm{\mathrm{P}}}}^{(1)}$ and ${\dot{\bm{\mathrm{P}}}}^{(2)}$ into $P$ overlapping patches respectively;
-10Divide each quaternion representation of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ and ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ into $P$ overlapping patches respectively;
-11Compute $QWSSIM$ values at the patch level using Eq. (17);
-12Obtain the refined focus map and fused patches at the patch level using Eqs. (19) and (20);
-Aggregate the patches of ${\{\dot{\bm{\mathrm{F}}}_{j}\}}_{j=1}^{P}$ to obtain the final result $\bm{\mathrm{F}}$;
-Algorithm 2 Quaternion Multi-focus Color Image Fusion
+To further enhance the quality of the final fused image, we propose a quaternion structural similarity refinement (QSSR) strategy.
+It refines the local patches of base-scale and detail-scale fusion results $\dot{\bm{\mathrm{F}}}_{1}$ and $\dot{\bm{\mathrm{F}}}_{2}$ using a patch-wise selection rule.
+The assumption is that the optimal fused patch should exhibit the highest similarity with the corresponding region in the source image that contains the most focused content.
+We first propose a weighted quaternion structure similarity ($WQ_{SSIM}$) measure.
+Weighted quaternion structural similarity measure.
+For each spatial location, let $\dot{\bm{\mathrm{f}}}$ denote a local patch of a base-scale or detail-scale fusion result $\dot{\bm{\mathrm{F}}}_{2}$ or $\dot{\bm{\mathrm{F}}}_{2}$.
+$\dot{\bm{\mathrm{p}}}_{1}$ and $\dot{\bm{\mathrm{p}}}_{2}$ represent the corresponding local patches extracted from $\dot{\bm{\mathrm{I}}}_{1}$ and $\dot{\bm{\mathrm{I}}}_{2}$ at the same spatial location.
+The $WQ_{SSIM}$ value is computed as follows:
+$WQ_{SSIM}({\dot{\bm{\mathrm{f}}}})={\tau}_{1}Q_{SSIM}(\dot{\bm{\mathrm{f}}},\dot{\bm{\mathrm{p}}}_{1})+{\tau}_{2}Q_{SSIM}(\dot{\bm{\mathrm{f}}},\dot{\bm{\mathrm{p}}}_{2})$ (20)
+where quaternion structural similarity ($Q_{SSIM}$) is defined in Definition 5 in Section II; ${\tau}_{1}$ and ${\tau}_{2}$ are the adaptive weights assigned to $\dot{\bm{\mathrm{p}}}_{1}$ and $\dot{\bm{\mathrm{p}}}_{2}$ respectively based on its detail-scale focus level.
+They are defined below:
+${\tau}_{1}=\frac{l_{D,1}}{l_{D,1}+l_{D,2}+\epsilon},\quad{\tau}_{2}=1-{\tau}_{1}$ (21)
+where $\epsilon$ is a small positive constant.
+$WQ_{SSIM}$ guides the selection of the optimal patch at each spatial location. The final fused patch $\dot{\bm{\mathrm{p}}}$ at each location is selected according to:
+$\dot{\bm{\mathrm{p}}}=\left\{\begin{aligned} &\dot{\bm{\mathrm{f}}}_{1},&{\rm if}\ WQ_{SSIM}({\dot{\bm{\mathrm{f}}}}_{1})>WQ_{SSIM}({\dot{\bm{\mathrm{f}}}}_{2})\\ &\dot{\bm{\mathrm{f}}}_{2},&{\rm if}\ WQ_{SSIM}({\dot{\bm{\mathrm{f}}}}_{1})\leq WQ_{SSIM}({\dot{\bm{\mathrm{f}}}}_{2})\end{aligned}\right$.
+(22)
+This strategy aims to effectively suppress artifacts and corrects misclassified regions in the fused outputs.
+Aggregating all selected patches according to Eq.
+(22) yields the quaternion fused result $\dot{\bm{\mathrm{F}}}$.
+Convert $\dot{\bm{\mathrm{F}}}$ back to the real domain to obtain the fused color image $\bm{\mathrm{F}}$.
 IV Experiments
-To evaluate the fusion performance of our framework, we conduct experiments on three public datasets.
+To evaluate the fusion performance of our QMCIF framework, we conduct experiments on three public datasets.
 Section A presents experimental settings like competing methods, datasets, and evaluation metrics.
-Sections B and C discuss the convergence and optimal parameters of the QSD model.
-To demonstrate the superiority of our method, Sections D and E present experimental results compared with other state-of-the-art approaches.
-Section F presents ablation studies on the mffw dataset [30] to verify the effectiveness of our QSD model, QBDF and QSSR strategies.
+Section B properties of the quaternion focal element decomposition (QFED) module that is a key component of the quaternion consistency-aware focus detection (QCAFD) method.
+To demonstrate the superiority of our QMCIF framework, Section C presents experimental results compared with state-of-the-art approaches.
+Section D presents ablation studies on the mffw dataset [29] to verify the effectiveness of our QCAFD module, QBDF and QSSR strategies.
 IV-A Experimental settings
-In our experiments, the patch size is set as $\sqrt{d}\times\sqrt{d}$ where $\sqrt{d}=8$ by default in the QSD model and base-scale fusion process.
-In the detail-scale fusion and the QSSR process, $\sqrt{d}$ is set to $5\times 10^{-3}MN$ [18], where $M$ and $N$ denote the height and width of the input image.
-The stopping criterion of QSD model is $51e-$ 5 in Algorithm 1.
+In our experiments, the patch size is set as $\sqrt{d}\times\sqrt{d}$ where $\sqrt{d}=8$ by default in the QCAFD method and base-scale fusion process.
+In the detail-scale fusion and QSSR process, $\sqrt{d}$ is set to $5\times 10^{-5}MN$ [30], where $M$ and $N$ denote the height and width of the input image.
+The stopping criterion of QFED is 1e51e-5 in Algorithm 1.
 All experiments are performed using Matlab 2016 on a workstation with a 2.90GHz Intel Core CPU and 16GB memory.
-Comparison methods.
-Nine representative multi-focus image fusion algorithms are selected as the comparison methods.
-They include sparse representation-based methods (SR) [10] as the baseline method, Guide-filter-based method (GFDF) [11], Multi-scale gradients-based method (MGIMF) [16], GAN-based method (MFIF-GAN) [31], Higher-order SVD-based method (QHOSVD) [19], unsupervised zero-shot based method (ZMFF) [32], Small-area-aware method (SAMF) [4], End-to-End based network (DBMFIF) [7], General-Image-Fusion-based (TCMOA) [33].
-SR belongs to channel-wise processing-based MCIF methods.
-GFDF, MGIMF, and SAMF are grayscale conversion-based MCIF methods.
-QHOSVD is a quaternion domain-based MCIF method.
-MFIF-GAN, ZMFF, DBMFIF, and TCMOA are deep-learning-based MCIF methods.
+Competing methods.
+Eleven representative multi-focus image fusion algorithms are selected as the comparison methods.
+They include a sparse coding-based (SR) method [14], a convolutional sparse coding-bsed (CSR) method [31], a multi-scale gradients-based (MGIMF) method [15], GAN-based (MFIF-GAN) method [6], a higher-order SVD-based (QHOSVD) method [21], unsupervised zero-shot based method (ZMFF) [32], a small-area-aware (SAMF) method [4], End-to-End based network (DBMFIF) [7], General-Image-Fusion-based (TCMOA) [33], a Swintransformer-based network (SwinMFF) [8] and a convolutional sparse representation-based unfolding network (MCCSR-Net) [34].
+SR and CSR are channel-wise processing-based methods.
+MGIMF and SAMF are grayscale conversion-based methods.
+QHOSVD is a quaternion domain-based method.
+MFIF-GAN, ZMFF, DBMFIF, TCMOA, SwinMFF and MCCSR-Net are deep-learning-based methods.
+TCMOA and SwinMFF are cross-channel attention-based methods.
 Datasets.
@@
 Metrics.
-Following [2], six objective metrics are adopted to evaluate the performance of two-color image fusion quantitatively.
+Following [1], six objective metrics are adopted to evaluate the performance of color image fusion quantitatively.
 They are the normalized mutual information $Q_{MI}$, gradient-based metric $Q_{G}$, phase congruency-based metric $Q_{P}$, structural similarity-based metric $Q_{E}$, $Q_{Y}$ and human perception-based metric $Q_{CB}$.
@@
 $Q_{CB}$ measures the preservation of contrast variations from source images in the fused image.
-IV-B Convergence Analysis
-The convergence analysis of the QSD model is provided in this subsection.
+IV-B Analysis of QFED
+This subsection analyzes our quaternion focal element decomposition (QFED) model that is a key component of QCAFD.
+We aim to validate QFED’s convergence, parameter sensitivity, and the impact of its each term on the whole fusion performance.
+Convergence Analysis.
 Fig.
-4 displays the evolution curves of the relative difference versus iterations of QSD model on the samples of lytro [34], mffw [30] and MFI-WHU datasets [35].
-The relative difference is calculated by the maximum value between $\|\dot{\bm{\mathrm{Z}}}{{}^{t+1}}-\dot{\bm{\mathrm{Z}}}{{}^{t}}\|{{}_{\infty}}$ and $\|\dot{\bm{\mathrm{D}}}{{}^{t+1}}-\dot{\bm{\mathrm{D}}}{{}^{t}}\|{{}_{\infty}}$.
+4 displays the evolution curves of the relative difference versus iterations of QFED on the samples of lytro [35], mffw [29] and MFI-WHU datasets [36].
+The relative difference is defined by the maximum value between $\|\dot{\bm{\mathrm{Z}}}{{}^{t+1}}-\dot{\bm{\mathrm{Z}}}{{}^{t}}\|{{}_{\infty}}$ and $\|\dot{\bm{\mathrm{D}}}{{}^{t+1}}-\dot{\bm{\mathrm{D}}}{{}^{t}}\|{{}_{\infty}}$.
 $\dot{\bm{\mathrm{Z}}}{{}^{t+1}}$ and $\dot{\bm{\mathrm{Z}}}{{}^{t}}$ denote the coefficient matrices at successive $t^{th}$ and ${(t+1)}^{th}$ iterations respectively.
 $\dot{\bm{\mathrm{D}}}{{}^{t+1}}$ and $\dot{\bm{\mathrm{D}}}{{}^{t}}$ denote the detail layers at successive $t^{th}$ and ${(t+1)}^{th}$ iterations respectively.
-The $\infty$ norm means the largest magnitude among each element of a vector.
-To improve the QSD speed, $\dot{\bm{\mathrm{B}}}{{}^{0}}$ is initialized as $\frac{\mathcal{F}(\dot{\bm{\mathrm{I}}})}{\mathcal{F}(\nabla_{1}^{T}\nabla_{1}+\nabla_{2}^{T}\nabla_{2})+1}$ and $\dot{\bm{\mathrm{D}}}{{}^{0}}$ is set to $\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}{{}^{0}}$ As displayed in Fig.
-4, QSD is convergent and its total number of iterations is 20.
-This can sufficiently guarantee convergence.
-Figure 4: Convergence analysis of QSD model on lytro, mffw and MFI-WHU datasets.
-IV-C Parameter Discussions
-Here, we discuss how three regularization parameters $\alpha$, $\beta$, and $\lambda$ affect the fusion performance of the QSD model.
-First, we show the visual effects of various values of $\alpha$, $\beta$ and $\lambda$.
-Then we comprehensively analyze and determine parameter settings.
-Figure 5: The fusion performance with different parameters on the lytro and the mffw datasets.
+The $\infty$ norm means the largest magnitude in each element of a vector.
+To improve the QFED speed, $\dot{\bm{\mathrm{B}}}{{}^{0}}$ is initialized as $\frac{\mathcal{F}(\dot{\bm{\mathrm{I}}})}{\mathcal{F}(\nabla_{1}^{T}\nabla_{1}+\nabla_{2}^{T}\nabla_{2})+1}$ and $\dot{\bm{\mathrm{D}}}{{}^{0}}$ is set to $\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}{{}^{0}}$ .
+As displayed in Fig.
+4, QFED is converges within 20 iterations.
+This demonstrates the computational efficiency and stability of the iterative solution of QFED.
+Figure 4: Convergence analysis of QFED model on lytro, mffw and MFI-WHU datasets.
+Parameter Sensitivity. We analyze the influence of regularization parameters $\alpha$, $\beta$, and $\lambda$ in the QFED model in Eq. (3).
+Figure 5: Fusion performance with different parameters on the lytro and the mffw datasets.
 (a) $Q_{E}+Q_{CB}$ results when parameters $\alpha$ and $\beta$ vary on the mffw dataset.
 (b) The $Q_{MI}$ value with respect to parameter $\lambda$ on both lytro and mffw datasets.
-In our QSD model, parameter $\alpha$ controls the base-scale layer while parameter $\beta$ handles the detail-scale layer.
-The final fused image not only preserves the most detailed parts of the inputs but also maintains the consistency of the background.
-To handle the three regularization parameters in our QSD model, we fix parameter $\lambda=0.05$ and adjust the others each time.
-$\alpha$ and $\beta$ were sequentially set values of [0.5, 1, 1.5, 2, 2.5, 3].
-Then $\alpha$ and $\beta$ were set to 1.5 and 2 and the variation values of $\lambda$ were sequentially set values of [0.05, 0.06, 0.07, 0.08, 0.09, 0.1].
-To effectively estimate the fusion performance, the combination index $Q_{E}+Q_{CB}$ approximates the fused images in terms of detail and base parts.
-Since $\lambda$ is not remarkably sensitive to the final fusion quality, index $Q_{MI}$ is used to quantitatively analyze the fused results.
-To quantitatively analyze the sensitivity of QSD versus these parameters, Fig.
-5 displays the $Q_{E}+Q_{CB}$ results of QSD with the samples of mffw dataset.
+In QFED, parameter $\alpha$ controls the base-scale layer while parameter $\beta$ handles the detail-scale layer.
+The final fused image not only preserves the most detailed regions of the inputs but also maintains the consistency of the background.
+To handle the three regularization parameters in our QFED model, we fix parameter $\lambda=0.05$ and adjust the others each time.
+$\alpha$ and $\beta$ are sequentially set values of [0.5, 1, 1.5, 2, 2.5, 3].
+Then $\alpha$ and $\beta$ are set to 1.5 and 2 and the variation values of $\lambda$ are sequentially set values of [0.05, 0.06, 0.07, 0.08, 0.09, 0.1].
+To effectively estimate the fusion performance, the combined quality metric $Q_{E}+Q_{CB}$ jointly assesses structural consistency and perceptual quality.
+Since $\lambda$ is not remarkably sensitive to the final fusion quality, $Q_{MI}$ is used to quantitatively analyze the fused results.
+We select the values corresponding to the maximum quality measure for the balance between detail preservation and background consistency.
+Fig.
+5 (a) displays the $Q_{E}+Q_{CB}$ results of QFED with the samples of mffw dataset.
+$\alpha$ and $\beta$ are set to 1.5 and 2 on mffw dataset as this setting yields the highest combined quality score.
+For lytro and MFI-WHU datasets, we select the optimal parameters $\alpha$ and $\beta$ in this manner and they are set to 1.5 and 0.5.
 In Fig.
-5 (a), $\alpha$ and $\beta$ are set to 1.5 and 2 on mffw dataset and $\beta$ are set to 0.5 on lytro and MFI-WHU datasets.
-$\lambda=5e-2$ is set for all datasets in Fig.
-5 (b).
-TABLE II: Performance comparison of multi-focus image fusion methods in two-color image fusion.
+5 (b), $Q_{MI}$ achieves its peak performance at $\lambda=0.05$.
+We adopt $\lambda=0.05$ consistently for all datasets since $\lambda$ exhibits less sensitivity to fusion quality variations.
+The parameter settings are uniformly employed in all subsequent experiments and evaluations.
+TABLE II: Ablation study on each term of QFED. The best results are bolded.
+Term settings $Q_{MI}\uparrow$ $Q_{G}\uparrow$ $Q_{P}\uparrow$ $Q_{E}\uparrow$ $Q_{CB}\uparrow$
+$\bm{\alpha=0}$ 1.0580 0.7359 0.7488 0.8240 0.7298
+$\bm{\beta=0}$ 1.0052 0.7338 0.7282 0.8279 0.6980
+$\bm{\lambda=0}$ 1.0140 0.7324 0.7432 0.8259 0.7127
+QFED 1.0685 0.7374 0.7832 0.8234 0.7508
+Figure 6: Two partially focused color image fusion on lytro, mffw and MFI-WHU datasets.
+Effectiveness of QFED.
+To verify the effectiveness of the QFED model, we perform the ablation study on each term of QFED.
+Specifically, three different settings of $\alpha=0$, $\beta=0$, $\lambda=0$ and $\alpha\neq 0$, $\beta\neq 0$, $\lambda\neq 0$ are particularly considered in this experiment.
+Table II gives the quantitative fusion results obtained by removing each term of QFED separately.
+As can be observed, each term can influence the fusion performance at different degrees.
+When $\beta$ is set to zero, we may obtain more focus-related features in the detail-scale layer $\dot{\bm{\mathrm{D}}}$.
+However, it fails to balance a tradeoff in the dual-scale focus map generation and subsequent fusion process.
+QFED provides the best quantitative results.
+IV-C Performance Comparison
+This subsection discusses the quantitative and qualitative comparison results of color image fusion.
+Quantitative evaluation.
+Table III provides the average performance of various competing algorithms on three public datasets.
+Six evaluation metrics are employed to assess the quality of the fused images, with higher values indicating better performance.
+In this paper, the highest evaluation scores are highlighted in bold.
+On the lytro dataset [35], our framework achieves the highest metric values compared to other competing methods across gradient-based, structural similarity-based, and human-perception-based evaluation metrics.
+This improvement is attributed to accurate focus detection capability of our QMCIF framework.
+However, our framework is worse than MCCSR-Net and SAMF on the metric of $Q_{MI}$.
+This is because our patch-wise fusion strategy tends to preserve less image intensity information during the quaternion fusion and refinement process compared to the pixel-wise fusion strategy.
+[…]
+TABLE III: Performance comparison of multi-focus image fusion methods.
 Dataset Methods $Q_{MI}\uparrow$ $Q_{G}\uparrow$ $Q_{P}\uparrow$ $Q_{E}\uparrow$ $Q_{Y}\uparrow$ $Q_{CB}\uparrow$
-lytro [34] SR [10] 1.0642 0.7445 0.8168 0.8797 0.9688 0.7756
-GFDF [11] 1.1194 0.7524 0.8335 0.8799 0.9826 0.8021
+lytro [35]
+SR [14] 1.0642 0.7445 0.8168 0.8797 0.9688 0.7756
+CSR [31] 1.0030 0.7353 0.8294 0.8794 0.9337 0.7612
 MGIMF [15] 1.1154 0.7428 0.8106 0.8419 0.9857 0.7911
@@
 TCMOA [33] 0.9926 0.7401 0.8188 0.8782 0.9574 0.7597
-Ours 1.1656 0.7613 0.8486 0.8813 0.9896 0.8056
-mffw [30] SR [10] 0.9181 0.6278 0.5500 0.8078 0.7844 0.6503
-GFDF [11] 1.0878 0.7323 0.7630 0.8242 0.9522 0.7426
+SwinMFF [8] 0.7689 0.7013 0.7773 0.8257 0.8867 0.6413
+MCCSR-Net [34] 1.1920 0.7619 0.8456 0.8772 0.9886 0.8084
+Ours 1.1656 0.7603 0.8472 0.8814 0.9896 0.8108
+mffw [29]
+SR [14] 0.7181 0.6278 0.5500 0.8078 0.9349 0.6503
+CSR [31] 0.9023 0.7065 0.6903 0.8222 0.8556 0.6861
 MGIMF [15] 1.0529 0.7309 0.7363 0.8096 0.9507 0.7338
@@
 TCMOA [33] 0.7545 0.6034 0.5307 0.7874 0.7681 0.6475
-Ours 1.0887 0.7338 0.7579 0.8275 0.9569 0.7450
-MFI-WHU [35] SR [10] 1.1329 0.7294 0.7952 0.8424 0.9850 0.8236
-GFDF [11] 1.1565 0.7337 0.8011 0.8416 0.9891 0.8312
+SwinMFF [8] 0.7259 0.6312 0.6336 0.7445 0.7285 0.6050
+MCCSR-Net [34] 0.8891 0.6925 0.6898 0.7903 0.8961 0.7233
+Ours 1.0887 0.7348 0.7642 0.8275 0.9579 0.7456
+MFI-WHU [36]
+SR [14] 1.1329 0.7294 0.7952 0.8424 0.9850 0.8236
+CSR [31] 0.9802 0.7157 0.7897 0.8423 0.9427 0.7827
 MGIMF [15] 1.1269 0.7151 0.7782 0.7973 0.9825 0.8157
@@
 TCMOA [33] 0.8635 0.6389 0.7026 0.7748 0.9139 0.7655
+SwinMFF [8] 0.6921 0.6776 0.7436 0.7906 0.8406 0.6599
+MCCSR-Net [34] 1.2011 0.7316 0.7996 0.8336 0.9807 0.8298
 Ours 1.1743 0.7361 0.8017 0.8433 0.9887 0.8318
-Figure 6: Two partially focused color image fusion on $lytro$, $mffw$ and $MFI-WHU$ datasets.
-IV-D Two color image fusion
-This subsection discusses the quantitative and qualitative results to assess fusion quality of two-color image fusion.
-Quantitative evaluation.
-Table II provides the average performance of various competing algorithms on three public datasets.
-Six evaluation metrics are employed to assess the quality of the fused images, with higher values indicating better performance.
-In this paper, the highest evaluation scores are highlighted in bold.
-On the lytro dataset [34], our framework achieves the highest metric values compared to the other competing methods across gradient-based, structural similarity-based, and human-perception-based evaluation metrics.
-This improvement is attributed to accurate focus detection capability of our method.
-However, our framework is slightly worse than SAMF on the metric of $Q_{MI}$.
-This is because our patch-wise fusion strategy tends to preserve less image intensity information during the quaternion fusion and refinement process compared to the pixel-wise fusion strategy.
-Since the images on the mffw dataset are affected by severe defocus spread effects and contain more challenging low-detail regions [30], all competing methods perform worse on this dataset compared to lytro and MFI-WHU datasets.
-Our method achieves the highest fusion performance on $Q_{G}$, and $Q_{E}$ metrics on the mffw dataset, highlighting its superior capability in focus detection and detail preservation.
-In the $Q_{MI}$ results, our framework slightly outperforms GFDF thanks to the effective artifact removal of our QSSR strategy.
-On the metrics of $Q_{Y}$ and $Q_{CB}$, our framework is worse only than MFIF-GAN.
-This demonstrates the advantages of fusion and refinement provided by the GAN-based MCIF algorithm in mitigating defocus spread effects [31].
-On the MFI-WHU dataset, our method achieves the best fusion performance in $Q_{G}$ and $Q_{E}$ metrics, underscoring its effectiveness in detecting focus regions from small areas, particularly when compared to SAMF.
-However, in the $Q_{Y}$ metric, our method performs slightly worse than GFDF and MFIF-GAN since both methods exhibit strong focus detection capabilities.
-Extensive experiments on various datasets demonstrate our framework significantly outperforms QHOSVD in terms of detail preservation and artifact removal under real-world challenging scenarios.
+MFI-WHU dataset contains challenging small-area blurred structures.
+Our framework achieves the best fusion performance in $Q_{G}$ and $Q_{E}$ metrics, underscoring its effectiveness in detecting focus regions from small areas, particularly when compared to SAMF.
+Compared with QHOSVD and MCCSR-Net, QMCIF clearly preserves more accurate structural details and achieves the highest perceptual quality $Q_{CB}$.
+However, in the $Q_{Y}$ metric, our framework performs slightly worse than GFDF and MFIF-GAN since both methods exhibit strong pixel-wise focus detection capabilities and our patch-wise fusion strategy tends to preserve less information from the original image in the boundary regions.
+Extensive experiments on various datasets demonstrate that our framework significantly outperforms QHOSVD in terms of detail preservation and artifact removal under real-world challenging scenarios.
 Visual evaluation.
-To evaluate the effectiveness of detail preservation and artifact removal, our method is compared with nine state-of-the-art methods.
-The red and yellow bounding boxes highlight areas with quality degradation.
-They are enlarged to better illustrate differences in the focused and defocused regions across the input images and output fused results.
-The first two rows of Fig.
-6 present the fused results of all competing methods on the lytro dataset.
-The red bounding boxes reveal that the results of SR, MGIMF, SAMF, and QHOSVD exhibit significant blurring along the edges.
-MGIMF shows spatial artifacts in the regions within the yellow bounding box.
-This is because the suboptimal decision maps are generated by multi-scale hand-designed filters.
-Our framework generates a pleasant fused result with clear objects and without any artifacts.
-The third and the fourth rows of Fig.
-6 illustrate the fused results of all competing methods on the mffw dataset.
-The red bounding boxes indicate that the outputs of SR, GFDF, and QHOSVD contain numerous blurred pixels.
-Similarly, the yellow bounding box highlights pseudo-edges and ghosts in the boundary regions of the MGIMF result.
-Furthermore, the outputs of SR, GFDF, MFIF-GAN, MGIMF, and QHOSVD suffer from varying degrees of color distortion and spatial artifacts in low-gradient regions.
-Our framework gives the most satisfactory fused result because our framework sufficiently handles the low-grafient regions and prevents visual artifacts for uncertain boundary areas.
-The last two rows of Fig.
-6 show the fused images produced by all comparative methods on the MFI-WHU dataset.
-The input images contain small and near blurred areas.
-This makes it difficult to estimate the focus levels and accurately identify focused regions.
-As can be seen in the red and yellow bounding boxes of the results, it is evident that SR, GFDF, MFIF-GAN, MGIMF, QHOSVD, SAMF and TCMOA produce unclear regions such as electric wires and poles.
-In contrast, our method demonstrates superior accuracy in detecting and preserving these focused regions.
-IV-E Multiple color image fusion
-To verify the superiority of our method in multiple-color image fusion, we conduct the experiment on the mffw3 dataset.
-Figure 7: Three partially focused color image fusion on $mffw3$ dataset.
-The source 1 image is a near-focus image.
-The source 2 data is a middle-focused image.
-The source 3 data is a far-focused image.
-Qualitative evaluation.
-Since TCMOA [33] does not provide a pre-trained model, only nine methods are included in this comparison.
-The deep learning-based methods fuse the multiple color images according to [36].
-As displayed in Fig.
-7, the experimental results show that these comparison methods perform poorly in the multiple-color image fusion process.
-GFDF, MGIMF, ZMFF and SAMF contain noticeable visual artifacts in the background areas.
-In the red bounding boxes of the QHOSVD and MGIMF results, significant blurring is observed, particularly in the areas of the wings.
-Similarly, in the yellow bounding box of the SAMF result, the fused image is obstructed by defocused pixels within the patterns.
-SR shows obvious artifacts and color distortion.
-This may be because SR fails to provide accurate focus level for focus detection and it is highly sensitive to large misregistration and color differences in the source images.
-Fig.
-7 indicates the superiority of our framework in terms of focus detection and artifact removal in foreground, background and boundary regions.
-IV-F Ablation study
-Effectiveness of QSD.
-To verify the advantages of the QSD model, we perform the ablation experiments on the mffw dataset (the first 10 pairs of samples) with two input color images.
-Different settings of parameters in Eq.
-(3) are provided, we investigate how the different terms of QSD influence the overall fusion performance.
-Specifically, three different settings of $\alpha=0$, $\beta=0$, $\lambda=0$ and $\alpha\neq 0$, $\beta\neq 0$, $\lambda\neq 0$ are particularly considered in this experiment.
-Table III gives the quantitative fusion results with different settings of QSD.
-As can be observed, the different settings of terms can influence the fusion performance at different degrees.
-The QSD model provides the best quantitative results.
-Therefore, this experiment further verifies the effectiveness of the $\alpha$ term, $\beta$ term and $\lambda$ term in Eq.
-(3).
-TABLE III: Quantitative evaluation on the different settings of the QSD model, the QBDF and the QSSR strategies.
+To assess the effectiveness of focus detection and artifact suppression, we compare our framework against five recent state-of-the-art fusion approaches.
+More visual comparison results are seen in the supplementary material.
+The top row of Fig.
+6 presents the fusion results on the lytro dataset.
+As observed in the red bounding boxes, QHOSVD and TCMOA suffer from noticeable edge blurring in the building structures while MCCSR-Net introduces spatial artifacts in the same regions.
+In contrast, our framework produces a visually pleasing fusion result with clear object boundaries and no observable artifacts.
+The middle row of Fig.
+6 shows the results on the mffw dataset.
+The blue bounding boxes reveal that QHOSVD, SAMF, and TCMOA generate blurred pixels near the object boundaries, and the DBMFIF output contains pseudo-edges and ghosting artifacts in low-contrast regions.
+Our framework delivers the most visually satisfactory fusion result, effectively addressing the challenges posed by low-gradient backgrounds and uncertain boundary transitions.
+The bottom row of Fig.
+6 displays fusion results on the MFI-WHU dataset.
+This scene contains small and near-blurred structures such as electric wires and poles, making focus estimation particularly challenging.
+The red bounding boxes clearly show that QHOSVD, SAMF, TCMOA, and MCCSR-Net fail to preserve these fine structures.
+In contrast, our framework accurately preserves these focused regions, demonstrating its superior capability in fine-grained focus detection and high-fidelity structure retention.
+Our framework can be easily extended to multiple color image fusion with architectural modifications.
+As an example, we extend our QMCIF framework to fuse three partially focused color images.
+By converting the input images into multiple quaternion representations, we obtain the detail layer and coefficient matrix for each input quaternion representation individually and generate patch-wise dual-scale focus maps using Eqs.
+(16) and (17).
+We select the patch with the most focus level of input images.
+Then we obtain dual-scale fusion results and the final fused result simply applying the QBDF and QSSR strategies.
+More details of multiple color image fusion are seen in the supplementary material.
+Figure 7: Visual comparison of three partially focused color image fusion on lytro3 dataset.
+Figure 8: Visual comparison of three partially focused color image fusion on mffw3 dataset.
+We conduct experiments on the lytro3 [37] and mffw3 [29] datasets.
+We compare our framework with three representative approaches that are designed for multiple image fusion.
+The deep learning-based baseline ZMFF is used for this comparison [32].
+Figs.
+7 and 8 presents the visual comparison results.
+More comparison results are included in the supplementary material.
+In Fig.
+7, QHOSVD produces visibly blurred shapes in the region marked by the red bounding box, especially around the device contour.
+ZMFF also suffers from color inconsistency in the same region, leading to degraded structure preservation and unnatural appearance.
+In Fig.
+8, ZMFF exhibits noticeable visual artifacts in the background, degrading perceptual quality.
+In contrast, our framework produces clean and structurally consistent results across all scenes.
+It effectively eliminates visual artifacts and maintains accurate focus representation in both foreground and background regions, as well as at structural boundaries.
+Our framework offers a simple yet robust solution for complex fusion tasks.
+IV-D Ablation study
+In this subsection, we conduct an ablation study with different settings of our QMCIF framework on mffw dataset [29].
+TABLE IV: Quantitative evaluation on different settings of our QMCIF framework.
 Settings $Q_{MI}\uparrow$ $Q_{G}\uparrow$ $Q_{P}\uparrow$ $Q_{E}\uparrow$ $Q_{CB}\uparrow$
-Parameter $\bm{\alpha=0}$ 1.0580 0.7359 0.7488 0.8240 0.7298
-$\bm{\beta=0}$ 1.0052 0.7338 0.7282 0.8279 0.6980
-$\bm{\lambda=0}$ 1.0140 0.7324 0.7432 0.8259 0.7127
-QSD 1.0685 0.7374 0.7832 0.8234 0.7508
-Patch size $\bm{5\times 5}$ 1.0503 0.7376 0.7277 0.8263 0.7115
+Patch size for QCAFD $\bm{5\times 5}$ 1.0503 0.7376 0.7277 0.8263 0.7115
 $\bm{10\times 10}$ 1.0401 0.7367 0.7567 0.8276 0.7276
 $\bm{8\times 8}$ 1.0685 0.7374 0.7832 0.8234 0.7508
-Fusion QBDF-Base-scale 0.9312 0.7021 0.7113 0.8099 0.6830
+Fusion for QBDF and QSSR QBDF-base-scale 0.9312 0.7021 0.7113 0.8099 0.6830
 QBDF-Detail-scale 1.0388 0.7357 0.7457 0.8255 0.7228
 QSSR 1.0685 0.7374 0.7832 0.8234 0.7508
-Weight $\tau$ Same 1.0690 0.7346 0.7556 0.8183 0.7340
+Weight $\tau$ of QSSR Same 1.0690 0.7346 0.7556 0.8183 0.7340
 Adaptive 1.0685 0.7374 0.7832 0.8234 0.7508
-Representation Real 0.9555 0.7082 0.6849 0.8043 0.6748
+Representation of QMCIF Real 0.9555 0.7082 0.6849 0.8043 0.6748
 Quaternion 1.0685 0.7374 0.7832 0.8234 0.7508
-Table III gives the quantitative fusion results with different patch sizes of the QSD model. As can be seen, the size $8\times 8$ provides the best quantitative results among five metrics.
+Impact of patch size in QCAFD.
+Table IV shows the quantitative analysis of different patch sizes within the QCAFD module.
+For QCAFD, we try different patch sizes to perform the QFED process and the patch-wise base-scale focus map generation for QBDF and QSSR strategies.
+The optimal performance is achieved with a patch size of $8\times 8$, demonstrating balanced sensitivity to both high-texture and low-gradient regions.
 Effectiveness of QBDF and QSSR.
-To verify the advantages of the QBDF and the QSSR strategies, we perform the ablation experiments on the mffw data.
+To verify the advantages of the QBDF and QSSR strategies, we perform the ablation experiments on the mffw data.
 This includes the first 10 pairs of partially focused color images under various exposures and obvious misalignment due to severe defocus spread effects.
 First, the coefficient matrix and the detail layer of quaternion representation of each input color image are extracted using Algorithm 1.
-Then, base-scale and detail-scale fused results are obtained using the focus measures from Algorithm 2.
+Then, base-scale and detail-scale fused results are obtained using the dual-scale focus measures.
 These results are subsequently combined to generate the final output using the QSSR strategy.
-Figure 8: Visual comparison of the background-focused input and various fused results including QBDF-Base-scale, QBDF-Detail-scale and QSSR results.
+Figure 9: Visual comparison of the background-focused input and various fused results including base-scale fusion, detail-scale fusion and the QSSR strategy.
 The first row represents the input and fused results.
 The second row represents the differences between the fused results and corresponding background-focused input.
+The importance of the QBDF and QSSR strategies is highlighted through qualitative and quantitative comparisons.
 Fig.
-8 provides visual comparison of the background-focused input color image and three fused results.
-As can be observed, QBDF-Detail-scale fusion introduces noticeable artifacts in the wall regions and creates unrealistic boundary lines within the yellow bounding box.
-In contrast, the QBDF base-scale fusion achieves smoother integration in low-texture areas but fails to preserve fine details in structured regions such as piano keyboards, resulting in oversmoothed edges.
-QSSR addresses these problems and significantly improves fusion performance in low-gradient, small-area focused, and boundary regions.
-The analysis of the QBDF and the QSSR strategies is also quantitatively presented in this ablation study.
-Table III gives the quantitative fusion results with different scales of measures.
-As can be observed, QSSR provides the best quantitative results among five metrics.
-To verify the robustness of our QSSR strategy, The weight $\tau$ of QSSR is discussed here.
-There are two different settings for $\tau$ in Eq.
-(17).
-The adaptive weight $\tau$ in Eq.
-(18) provides the best quantitative results among five metrics.
+9 visually demonstrates while detail-scale fusion alone leads to artifacts and base-scale fusion smooths out fine details excessively.
+The QSSR strategy effectively preserves important structural details and eliminates artifacts, especially in boundary and low-gradient regions.
+The quantitative results in Table IV are align with this visual assessment.
+This indicates the improvements across multiple metrics when employing QSSR.
+Adaptive weighting in QSSR.
+We further evaluate the robustness of adaptive weighting in the QSSR strategy (Eq.
+(21)).
+Table IV reveals that adaptive weighting consistently outperforms fixed weighting across all evaluated metrics.
+This confirms that the adaptive weighting accurately assesses local structural similarity to optimize patch selection.
 Effectiveness of QMCIF.
-Table III also evaluates the impact of different representation strategies on fusion performance by comparing our QMCIF framework with conventional real-domain fusion method.
-To separate the benefits of quaternion representation, we replicate the fusion process in the real domain by converting color images to grayscale.
-Quantitative results demonstrate that the quaternion representation achieves superior performance across all metrics, particularly outpeforming real domain-based framework by $12\%$ in $Q_{CB}$.
+Table IV compares the quaternion representation with an equivalent real-domain under our framework.
+We replicate the fusion process in the real domain by converting color images to grayscale.
+Quantitative results in Table IV highlight inherent advantages of quaternion representation in preserving inter-channel correlations and structural coherence during fusion.
 V Conclusion
-We proposed a quaternion multi-focus color image fusion framework to perform high-quality color image fusion completely in the quaternion domain.
+This paper proposed a quaternion multi-focus color image fusion framework to perform high-quality color image fusion completely in the quaternion domain.
 This framework is flexible to be extended from fusing two color images at a time into fusing multiple color images simultaneously under complex challenging scenarios.
-This framework proposed a quaternion sparse decomposition model for high-precision focus detection of color images.
-This model jointly learns a low-rank structured coefficient matrix and a detail-scale layer.
-They both were utilized to estimate the focus level of the image.
-To solve the optimization problem of the QSD model, we developed an iterative algorithm under the framework of the quaternion alternating direction methods of multipliers.
+This framework proposed a quaternion consistency-aware focus detection method.
+It consists of a quaternion focal element decomposition (QFED) module that jointly learns a low-rank structured coefficient matrix and a detail-scale layer and a dual-scale focus map generation strategy for robust focus detection.
+To solve the optimization problem of the QFED model, we developed an iterative algorithm under the framework of the quaternion alternating direction methods of multipliers.
 A quaternion base-detail fusion strategy was also introduced for high-quality base-scale and detail-scale color image fusion individually.
 A quaternion structural similarity strategy was further introduced to balance the trade-off between focus information preservation and artifact removal.
-Our framework fuses two-scale quaternion matrices and compares them with the input quaternion representations to adaptively select the focused patches with a higher QWSSIM value to form the final fused results.
+Our framework fuses two-scale quaternion matrices and compares them with the input quaternion representations to adaptively select the focused patches with a higher $WQ_{SSIM}$ value to form the final fused results.
 Extensive experiment results have shown that our framework outperforms the state-of-the-art methods in terms of focus information preservation and artifact removal under real-world challenging scenarios.